### v2

# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

Abstract.
The celebrated Robinson–Schensted algorithm and each of its variants that have attracted substantial attention can be constructed using Fomin’s “growth diagram” construction from a modular lattice that is also a weighted-differential poset. In this paper we classify all such “Fomin” lattices that meet certain criteria; the main criterion is that the lattice is distributive. Intuitively, these criteria seem excessively strict, but all known Fomin lattices satisfy all of these criteria, with the sole exception of one family that is not even distributive, the Young–Fibonacci lattices and cartesian products involving them. We discover a new class of Fomin lattices, but unfortunately they cannot be used to construct Robinson–Schensted algorithms.
Key words and phrases:
differential poset, Fomin lattice, graded graph, growth diagram, Robinson–Schensted algorithm, Young diagram, Young tableaux
2020 Mathematics Subject Classification:
Primary 06A11; Secondary 05A17, 06B99
Note added to the 2nd version: Paragraphs 6.5, 6.6, and 6.7 of the 1st version were incorrect and have been replaced. Their conclusions remain unchanged. Also, a new class of Fomin lattices has been added as item (11) of para. 2.3: the “vertical” joining of an arbitrary doubly-infinite sequence boolean algebras $B_{1}$ or $B_{2}$.
1. Introduction
The celebrated Robinson–Schensted algorithm[Rob1938a][Schen1961a][Fom1994a][Fom1995a] and each of its variants that have attracted substantial attention can be constructed[Wor2023c]*sec. 2 using Fomin’s “growth diagram” construction from a modular lattice that is also a (weighted-)differential poset[Stan1988a]. We define these requirements as a “Fomin lattice” that has a minimum element $\hat{0}$. In this paper we classify all Fomin lattices that meet certain criteria. Our main criterion is that the lattice is distributive, with secondary criteria on the set of points (the join-irreducible elements of the lattice). Intuitively, these criteria seem excessively strict, but all known Fomin lattices satisfy all of these criteria, with the sole exception of one family that is not even distributive, the Young–Fibonacci lattices. This work adds to the classes of lattices within which all of the Fomin lattices have been classified.
We find a class of Fomin lattices that appear to be new to the literature: Let $L$ be the “vertical” joining of an arbitrary doubly-infinite sequence of 2-chains or $B_{1}$ (boolean algebras on 1 element), and diamonds or $B_{2}$ (boolean algebras on 2 elements). Then $L$ with the weighting is Fomin, but unfortunately has no $\hat{0}$. See an example in fig. 4 and para. 2(11).
The history of this problem in the literature is discussed in para. 2. Para. 2 enumerates the relevant properties of structures. All known Fomin lattices are described in para. 2. They consist of Young’s lattice of partitions, the shifted Young’s lattice of strict partitions, the lattices of partitions into a limited number of parts, the Young-Fibonacci lattices, the cylindric partition lattices, the lattice of ideals of a “skew strip”, the new class of lattices, some trivial cases, and certain derivatives of other Fomin lattices. The previous classification theorems for Fomin lattices are discussed in para. 2.
Our new classification theorem is in sec. 7. The class of lattices classified by our theorem is incomparable with the classes of previous classification theorems; it does not supersede them. However, it is an incremental contribution to the classification of Fomin lattices (and thus Robinson–Schensted algorithms) and is the only one that applies to “weighted-differential” lattices, that is, where the weight on coverings (or points) is not identically 1.
Our proof techniques are elementary. They consist of a careful accounting of the consequences of the differential condition (7) in lattices that satisfy our criteria. This “local characterization” (sec. 5) includes a succession of restatements equivalent to the differential condition (lem. 5.7, 5.17, 5.18, 5.22, 5.23, 5.27, 5.28, and 5.29). These results culminate in lem. 5.30 and 5.31, that no point covers three distinct points or is covered by three distinct points. We then proceed to the “global characterization” (sec. 6) of how the points can be placed on a two-dimensional grid. This culminates in the classification theorem, th. 7.1.
Finally, we discuss possible future directions for research (sec. 8), which assesses the prospect of various relaxations of the criteria of the classification theorem.
Thus this article consists of:
  • background information including the previous literature, a catalog of properties, known Fomin lattices, and a recapitulation of properties of the Cartesian product (sec. 2),
  • a statement of the definitions essential for the classification (sec. 3),
  • a construction supporting the discussion of future directions (sec. 4),
  • the characterization of Fomin lattices that meet our criteria (sec. 5 and 6),
  • the classification theorem for Fomin lattices that meet our criteria (sec. 7), and
  • a discussion of future directions for research (sec. 8).
The development essential to prove our classification theorem 7.1 is contained in sec. 3, 5, 6, and 7.
2. Background
Previous literature
The literature on differential structures can be classified based on the generality of the structures being considered. The generality can be described along two dimensions:
  1. (1)
    the generality of the underlying structures (lattice, poset, etc.) being considered and
  2. (2)
    whether the covering pairs of elements of the structure are all given the weight 1 (the unweighted case) or are allowed varying weights (the weighted case).
In regard to the underlying structures, from most general to least general, the major categories are
  1. (1)
    dual graded graph (two graphs or multi-graphs on the same set of vertices with the same grading with separate weights or edge multiplicities in the two graphs between pairs of points),
  2. (2)
    graded poset,
  3. (3)
    lattice (which by the differential property must be modular),
  4. (4)
    distributive lattice, and
  5. (5)
    distributive lattice with the further conditions of th. 7.1.
Significant works in the literature, and distinctive parts of them, are classified in the following table. In some cases a work’s definitions and general analysis are considerably more general than all of the examples described in the work; in that case the classification of the definitions are marked “D” and classification of the examples are marked “E”. Some works that are largely a description of examples (often novel examples) are only marked with “E”. The works with the best previous classification results (detailed at the end of para. 2) are listed with “C” in the appropriate category.
Weighting
Weighted Unweighted
Underlying structure Dual graded graph Fomin [Fom1994a]*Def. 1.3.4 Qing [Qing2008a]*sec. 4.2 E
Graded poset Stanley [Stan1988a]*Def. 1.1 DC
Lewis [Lew2007a]*sec. 3 E
Lattice (modular) Stanley [Stan1990a]*sec. 3
this article D
Stanley [Stan1988a] E
Byrnes [Byrn2012a]*Ch. 6 C
Distributive lattice Fomin [Fom1994a]*sec. 2.2 Elizalde [El2025a] E
Distributive lattice with further restrictions this article E
Properties considered in this article
A large number of properties of structures are considered in this area of combinatorics. This section catalogs the properties that are relevant to this article and their definitions. This article classifies in sec. 7 structures which have all of these properties. The properties are here generally ordered from weakest to strongest. Some theorems relating the properties are included. For a casual reading of this article, this section can be skipped and referred back to as necessary. Section 3 gives terse definitions that suffice for reading characterizations in this article.
  1. (A)
    The structure is a dual graded graph (DGG): [Fom1994a]*Def. 1.3.4 There is a graded set of vertices. In the original formulation each grade contains a finite number of vertices, but we do not include that criterion here. (But see property (D).) Vertices in a grade have edges to some vertices in adjacent grades, forming a graph. Each edge has an assigned weight, with the weights being drawn from a set of values $V$. In the original formulation, each edge is an otherwise unlabeled multi-edge, so the weights are positive integers ($V=\mathbb{Z}$), but we do not include that criterion here.
    A dual graded graph is two such graphs on the same graded set of vertices, and thus there are up to two weights between any pair of vertices in adjacent grades. The edges of the first graph are conceptualized as being directed upward (toward higher grades) and the edges of the second graph are conceptualized as being directed downward.
    There is some ambiguity in the presentation of a DGG; if a pair of vertices has a non-zero weight between them in only one graph, we may add an edge between them in the other graph that has a weight of 0 without changing any of the essential properties of the DGG. This allows the two graphs to have the same set of edges, with the two graphs differing only in the weights of the edges.
    Critically, the DGG has the differential property: It has a constant $r$ called the differential degree and the DGG is called $r$-differential. Assume that weight 0 edges have been added so that both graphs have the same edges. Let $x\lessdot y$ mean that $y$ is in the grade above $x$ and $(x,y)$ is an edge. When $x\lessdot y$, let $w_{1}(x\lessdot y)$ be the weight of the edge $(x,y)$ in the first graph and $w_{2}(x\lessdot y)$ be the weight of the edge $(x,y)$ in the second graph. Then the differential property is that for every pair of vertices $x$ and $y$ in the same grade,
    $\sum_{z\,|\,x\lessdot z,y\lessdot z}w_{1}(x\lessdot z)w_{2}(x\lessdot z)-\sum_{v\,|\,v\lessdot x,v\lessdot y}w_{2}(v\lessdot x)w_{1}(v\lessdot y)=\begin{cases}0&\text{if }x\neq y\\ r&\text{if }x=y\end{cases}$ (1)
  2. (B)
    The structure is a differential poset: [Stan1988a] This is a narrower type of DGG. Both graphs have the same set of edges and the weights on the edges of the second graph are all 1. In the original formulation, the weights on the edges of the first graph are also 1, but we do not include that criterion here. (But see property (H).) Differential posets with weights that are not 1 are also called weighted-differential posets.
    The graphs are the Hasse diagram of a poset. In the original formulation, the poset is assumed to have a minimum element, $\hat{0}$, but we do not include that criterion here. A differential poset is required to be unique-cover-modular.
Definition 2.1.
A poset $P$ is cover-modular [^1] We choose this name because of its resemblance to the property of coverings in modular lattices. But in general, $P$ is not necessarily modular, indeed, it need not even be a lattice. A poset with this property and finite length is called a “modular poset” in [Birk1967a]*§II.8. if when $x,y\in P$ and $x$ and $y$ both cover an element of $P$, then they are both covered by an element of $P$ (which may not be unique). Dually, we require that when $x,y\in P$ and $x$ and $y$ both are covered by an element of $P$, then they both cover an element of $P$ (which may not be unique).
Definition 2.2.
A poset $P$ is unique-cover-modular [^2] This property is called a “uniquely modular poset” in [Proct1982a]*sec. 5. if when $x,y\in P$ and either $x$ and $y$ both cover an element of $P$ or $x$ and $y$ are covered by an element of $P$, then they both cover a unique element of $P$ and they are both covered by a unique element of $P$.
Theorem 2.3.
[Birk1967a]*§II.8 Th. 14 If a locally-finite poset has $\hat{0}$ and is cover-modular, then it is graded.
  1. (C)
    The structure is locally-finite: In both the DGG and differential poset formulations, the structure (seen as a poset) is assumed to be locally-finite.
  2. (D)
    The structure has finite covers: In both the DGG and differential poset formulations, every vertex has a finite number of upward and downward edges, or a finite number of upward and downward covers.
    [Fom1994a]*sec. 1.1 requires each grade of vertices to be finite. If the structure has a $\hat{0}$, this is equivalent to finite covers, but without a $\hat{0}$, it is a stronger property. (See also property (H).)
  3. (E)
    The structure has unique-modular-covers and the weights are projective-constant: In this case, if $x\neq y$, (1) is satisfied automatically. If we define $w(x\lessdot y)=w_{1}(x\lessdot y)w_{2}(x\lessdot y)$, then the $x=y$ case of (1) is equivalent to: For every vertex $x$,
    $\sum_{z\,|\,x\lessdot z}w(x\lessdot z)-\sum_{v\,|\,v\lessdot x}w(v\lessdot x)=r$ (2)
    Thus if we require unique-modular-covering and projective-constancy, the DGG and weighted-differential poset structures effectively coincide with only one weight function. Differential posets that are not weighted-differential are automatically projective-constant.
Definition 2.4.
The weights on a DGG or differential poset are projective-constant iff (for a DGG) the weight functions $w_{1}$ and $w_{2}$ or (for a differential poset) the weight function $w$ have the following property: the function $f$ is equal on coverings which are cover-projective: If $x\vee y$ covers $x$ and $y$ and $x\wedge y$ is covered by $x$ and $y$, then $f(x\wedge y\lessdot x)=f(y\lessdot x\vee y)$ and $f(x\wedge y\lessdot y)=f(x\lessdot x\vee y)$.
  1. (F)
    The structure is a differential poset that is a lattice.
Lemma 2.5.
[Birk1967a]*§II.8 Th. 16 If a lattice is locally-finite and cover-modular, then it is modular.
  1. (G)
    The structure is a differential poset that is a modular lattice: If the structure is a locally-finite modular lattice, then if the weight function is projective-constant, it is constant on projectivity classes[Birk1967a]*§I.7 Def. of coverings.
We define “Fomin lattices” to be structures with the above set of properties.
  1. (H)
    The structure has a $\hat{0}$: We consider the existence of a minimum vertex as a separate property. [Stan1988a] defines all differential posets to have a $\hat{0}$.
  2. (I)
    The weighting and differential degree are positive.
Definition 2.6.
A weighting $w$ and differential degree $r$ (taken together) are positive if the set of values $V$ (the target set of $w$) is $\mathbb{Z}$, all of the values of $w$ are positive, and $r$ is positive.
Depending on the degree of abstraction of the discussion, the requirements for the set of values $V$ varies. The minimum requirement is that $V$ be a module over a unitial ring of scalars $S$. Usually we will be unconcerned with the details of $S$ other than it necessarily contains a homomorphic image of $\mathbb{Z}$ and thus we can multiply elements of $V$ by integer coefficients. If $V=\mathbb{Z}$, then $\mathbb{Z}$ is isomorphically embedded in $S$.
Lemma 2.7.
If $L$ is a Fomin lattice with $\hat{0}$, $L$ is not the trivial (one-element) lattice, $V=\mathbb{Z}$, and all values of $w$ are positive, then $r$ is positive.
Proof.
For $x=\hat{0}$, (2) becomes $r=\sum_{z\,|\,x\lessdot z}w(\hat{0}\lessdot z)$. But by hypothesis, there is at least one $z$ in the range of the sum and $w(\hat{0}\lessdot z)>0$, which ensures the sum is positive. That proves $r$ is positive. ∎
In light of the above lemma, we will abuse language by using “$w$ is positive” or “the lattice is positive” to mean that the weighting and differential degree are positive, that is, $V$ is $\mathbb{Z}$, $r>0$, and all of the values of $w$ are positive.
  1. (J)
    The structure is a distributive lattice: In this case, we call the lattice $L$ and define $P$ to be the poset of points (join-irreducibles) of $L$. If the lattice is finitary, it is isomorphic to the set of finite order ideals of $P$.
  2. (K)
    The structure is a lattice and the poset of points $P$ is unique-cover-modular: Note that this requirement is on the poset of points, not a requirement on the structure itself (which all Fomin lattices satisfy).
  3. (L)
    The structure is a lattice that cannot be factored into two nontrivial lattices: If the lattice is distributive this is equivalent to that $P$ cannot be partitioned into two disjoint non-empty posets $P=P_{1}\sqcup P_{2}$. [^3] We use $\sqcup$ for disjoint unions of sets and posets because “disjoint union” is the coproduct in the categories of sets and posets.
Known Fomin lattices
All known Fomin lattices can be constructed as follows:
  1. (1)
    Any lattice satisfying property (A) can be made into a degenerate Fomin lattice by setting $r=0$ and $w=0$.
  2. (2)
    [Fom1994a]*Exam. 2.1.2 and 2.2.7 Young’s lattice $\mathbb{Y}$, the lattice of partitions, with $r=1$ and $w=1$ is a positive Fomin lattice with $\hat{0}$. See fig. 1. (There are additional weightings $w$ on this lattice that make it Fomin but are not positive.)
    (a) Lattice $\mathbb{Y}$ with weights of the coverings
    (b) The quadrant $J(\mathbb{Y})$ with weights of the points
    Figure 1. Young’s lattice $\mathbb{Y}$, the lattice of partitions, with its canonical weighting.
  3. (3)
    [Sag1979b][Wor1984a][Fom1994a]*Exam. 2.2.8 The shifted Young’s Lattice $\mathbb{SY}$, the lattice of partitions into unequal parts, with $r=1$ and $w=1\textup{ or }2$ is a positive Fomin lattice with $\hat{0}$. See fig. 2.
    (a) Lattice $\mathbb{SY}$ with weights of the coverings
    (b) The octant $J(\mathbb{SY})$ with weights of the points
    Figure 2. The shifted Young’s lattice $\mathbb{SY}$, the lattice of strict partitions, with its canonical weighting.
  4. (4)
    [Fom1994a]*Exam. 2.2.5 For any integer $k\geq 1$, the $k$-row Young’s lattice, $\mathbb{Y}_{k}$, the lattice of partitions into at most $k$ parts (or dually, into an arbitrary number of parts that are all $\leq k$), with $r=k$ and $w$ having a particular range of values is a positive Fomin lattice with $\hat{0}$. See $\mathbb{Y}_{3}$ in fig. 3.
    (a) Lattice $\mathbb{Y}_{3}$ with weights of the coverings
    (b) $J(\mathbb{Y}_{3})$ with weights of the points
    Figure 3. The lattice $\mathbb{Y}_{3}$ of partitions with $\leq 3$ rows, with its canonical weighting.
  5. (5)
    [Fom1994a]*Exam. 2.2.1 The upward semi-infinite chain $\mathbb{N}$ with $r=k$ and $w(i)=k(i+1)$ (which is isomorphic to $\mathbb{Y}_{1}$, a special case of (4)) is a positive Fomin lattice with $\hat{0}$.
  6. (6)
    [Fom1994a]*Exam. 2.2.9 The infinite chain $\mathbb{Z}$ with and $w(i)=ri$ is a Fomin lattice but has no $\hat{0}$ and is not positive.
  7. (7)
    [Stan1988a]*sec. 5[Roby1991a]*Ch. 5 For any integer $k\geq 1$, the Young–Fibonacci lattice of degree $k$, $\mathbb{YF}_{k}$, with $r=k$ and $w=1$ is a positive Fomin lattice with $\hat{0}$.
  8. (8)
    [El2025a]*sec. 5 For any integers $d,L\geq 1$, the cylindric partition lattice $\mathbb{CP}_{dL}$ with $r=0$ and $w=1$ is a Fomin lattice but is not positive and has no $\hat{0}$.444[El2025a] uses the notation $\mathcal{L}_{d,L}$ for $\mathbb{CP}_{dL}$. The cylindric partitions have been known since at least [GessKratt1997a] but [El2025a] seems to be the first report that all of the $\mathbb{CP}_{dL}$ are 0-differential. See (9): [Fom1994a] reports that the $\mathbb{CP}_{2t}$ are 0-differential.
  9. (9)
    [Fom1994a]*Exam. 2.2.10 defines $t\textit{-SkewStrip}$, which is a special case of (8): $t\textit{-SkewStrip}=\mathbb{CP}_{2t}$.
  10. (10)
    [Fom1994a]*Exam. 2.2.11 $\mathbb{SS}_{t}$ is the lattice of non-trivial ideals of $t\textit{-SkewStrip}=\mathbb{CP}_{2t}$, those that are neither empty nor all of $t\textit{-SkewStrip}$, with $r=0$ and $w=1\textup{ or }2$ is a Fomin lattice but is not positive and has no $\hat{0}$.
  11. (11)
    Let $L$ be the “vertical” joining of an arbitrary doubly-infinite sequence of 2-chains or $B_{1}$ (boolean algebras on 1 element), and diamonds or $B_{2}$ (boolean algebras on 2 elements). Then $L$ with $r=0$ and $w=1\textup{ or }2$ is a Fomin lattice but is not positive and has no $\hat{0}$. See an example in fig. 4.
    (a) A section of the poset of points, with the weights of the points
    (b) A section of the lattice, with the weights of the coverings, labeled with the generators of the ideals
    Figure 4. Example of a Fomin lattice of type (11)
  12. (12)
    [Fom1994a]*Lem. 2.2.3 Any cartesian product of a finite number of Fomin lattices that all have the same value set $V$, with the weight of a covering in the product lattice equal to the weight within the factor from which the covering is derived and the differential degree equal to the sum of the differential degrees of the factors, is a Fomin lattice. The product lattice has $\hat{0}$ if all factors have $\hat{0}$ and is positive if all factors are positive.
  13. (13)
    Given a Fomin lattice, the same lattice but multiplying the differential degree and weighting by the same constant gives a Fomin lattice. The resulting lattice is positive if the original lattice is positive and the constant is a positive integer.
Previous classifications
The previous classifications of Fomin lattices with $\hat{0}$ are:
  1. (1)
    [Stan1988a]*Prop. 5.5 If a Fomin lattice with $\hat{0}$ is distributive and has $w=1$, then it is isomorphic to $\mathbb{Y}^{\times r}$. [^5] We use $\bullet^{\times n}$ to denote the $n$-fold cartesian exponential to avoid ambiguity with other uses of superscripts.
  2. (2)
    [Byrn2012a]*Th. 1.3 If a Fomin lattice with $\hat{0}$ has $r=1$ and $w=1$, then it is either $\mathbb{Y}$ or $\mathbb{YF}_{1}$.
In particular, there are no previous classifications for the weighted case, when $w$ is not identically $1$.
Cartesian product
Para. 2 item (12) above shows that cartesian product preserves Fomin lattices. Surprisingly, the converse is also true.
Theorem 2.8.
If $L$ is a Fomin lattice with differential degree $r$ and weighting $w$, and $L$ is the cartesian product of two lattices $L=L_{1}\times L_{2}$, then
  1. (1)
    $L_{1}$ is a Fomin lattice with some differential degree $r_{1}$ and weighting $w_{1}$.
  2. (2)
    $L_{2}$ is a Fomin lattice with some differential degree $r_{2}$ and weighting $w_{2}$.
  3. (3)
    $r=r_{1}+r_{2}$.
  4. (4)
    For any $x\lessdot y$ in $L_{1}$ and $z$ in $L_{2}$, $w((x,z)\lessdot(y,z))=w_{1}(x\lessdot y)$.
  5. (5)
    For any $w$ in $L_{1}$ and $x\lessdot y$ in $L_{2}$, $w((v,x)\lessdot(v,y))=w_{2}(x\lessdot y)$.
Proof.
First we show that for any $x\lessdot y$ in $L_{1}$ and any $z_{1},z_{2}\in L_{2}$,
$w((x,z_{1})\lessdot(y,z_{1}))$ $=w((x,z_{2})\lessdot(y,z_{2})).$ (3)
Because of property (E), $w$ is constant on cover-projective pairs of elements of $L$, so equation (3) holds if $z_{1}\lessdot z_{2}$. Since any two elements $z_{1},z_{2}\in L_{2}$ can be connected by a finite chain of upward and downward covers, equation (3) holds for any $z_{1}$ and $z_{2}$. Thus the $w_{1}$ in conclusion (4) exists.
Similarly, for any $w_{1},w_{2}\in L_{1}$ and any $x\lessdot y$ in $L_{2}$
$w((w_{1},x)\lessdot(w_{1},y))$ $=w((w_{2},x)\lessdot(w_{2},y)).$ (4)
Thus the $w_{2}$ in conclusion (5) exists.
We now show that $r_{1}$ and $r_{2}$ exist, that is, that $L_{1}$ is a Fomin lattice with differential degree $r_{1}$ and weighting $w_{1}$ and similarly for $L_{2}$. Consider any $x\in L_{1}$ and any $z_{1},z_{2}\in L_{2}$. Equation (2) for $(x,z_{1})$ is:
$\sum_{(p,q)\,|\,(p,q)\lessdot(x,z_{1})}w((p,q)\lessdot(x,z_{1}))+r$ $=\sum_{(t,u)\,|\,(x,z_{1})\lessdot(t,u)}w((x,z_{1})\lessdot(t,u))$
$\sum_{p\,|\,p\lessdot x}w((p,z_{1})\lessdot(x,z_{1}))+\sum_{q\,|\,q\lessdot z_{1}}w((x,q)\lessdot(x,z_{1}))+r$ $=\sum_{t\,|\,x\lessdot t}w((x,z_{1})\lessdot(t,z_{1}))+\sum_{u\,|\,z_{1}\lessdot u}w((x,z_{1})\lessdot(x,u))$
$\sum_{p\,|\,p\lessdot x}w_{1}(p\lessdot x)+\sum_{q\,|\,q\lessdot z_{1}}w_{2}(q\lessdot z_{1})+r$ $=\sum_{t\,|\,x\lessdot t}w_{1}(x\lessdot t)+\sum_{u\,|\,z_{1}\lessdot u}w_{2}(z_{1}\lessdot u)$ (5)
Similarly, equation (2) for $(x,z_{2})$ is:
$\sum_{p\,|\,p\lessdot x}w_{1}(p\lessdot x)+\sum_{q\,|\,q\lessdot z_{2}}w_{2}(q\lessdot z_{2})+r$ $=\sum_{t\,|\,x\lessdot t}w_{1}(x\lessdot t)+\sum_{u\,|\,z_{2}\lessdot u}w_{2}(z_{2}\lessdot u)$ (6)
Subtracting (6) from (5) gives
$\sum_{q\,|\,q\lessdot z_{1}}w_{2}(q\lessdot z_{1})-\sum_{q\,|\,q\lessdot z_{2}}w_{2}(q\lessdot z_{2})$ $=\sum_{u\,|\,z_{1}\lessdot u}w_{2}(z_{1}\lessdot u)-\sum_{u\,|\,z_{2}\lessdot u}w_{2}(z_{2}\lessdot u)$
$\sum_{u\,|\,z_{1}\lessdot u}w_{2}(z_{1}\lessdot u)-\sum_{q\,|\,q\lessdot z_{1}}w_{2}(q\lessdot z_{1})$ $=\sum_{u\,|\,z_{2}\lessdot u}w_{2}(z_{2}\lessdot u)-\sum_{q\,|\,q\lessdot z_{2}}w_{2}(q\lessdot z_{2})$
This shows that $\sum_{u\,|\,z\lessdot u}w_{2}(z\lessdot u)-\sum_{q\,|\,q\lessdot z}w_{2}(q\lessdot z)$ is independent of $z\in L_{1}$. We define $r_{2}$ as this value.
Similarly, we show that $\sum_{t\,|\,z\lessdot t}w_{1}(z\lessdot t)-\sum_{q\,|\,q\lessdot z}w_{1}(q\lessdot z)$ is independent of $z\in L_{1}$, and we define $r_{1}$ as this value.
From these facts it is straightforward to prove the conclusions of the theorem. ∎
Lemma 2.9.
If $L$ is a Fomin lattice with differential degree $r$ and weighting $w$, then
  1. (1)
    If $L$ has a minimum element $\hat{0}_{L}$, then $L_{1}$ has a minimum element $\hat{0}_{L_{1}}$ and $L_{2}$ has a minimum element $\hat{0}_{L_{2}}$.
  2. (2)
    If $L$ has a minimum element $\hat{0}_{L}$ and is positive, and $L_{1}$ and $L_{2}$ are not the trivial lattice, then $L_{1}$ and $L_{2}$ are positive.
  3. (3)
    If $L$ is distributive, then $L_{1}$ and $L_{2}$ are distributive.
Proof.
Regarding (1) and (3): These are straightforward lattice reasoning.
Regarding (2): By the construction in th. 2.8, $L_{1}$ has some differential degree $r_{1}$ and weighting $w_{1}$ and $L_{2}$ has differential degree $r_{2}$ and weighting $w_{2}$.; Every value of $w_{1}$ and every value of $w_{2}$ are values of $w$ for suitable arguments, so all values of $w_{1}$ and $w_{2}$ are elements of the set of values $V$ of $L$. Thus the set of values $V_{1}$ and $V_{2}$ of $L_{1}$ and $L_{2}$ can be considered to be $\mathbb{Z}$, and all values of $w_{1}$ and $w_{2}$ are positive. Since $L$ has a $\hat{0}_{L}$, $L_{1}$ has a $\hat{0}_{L_{1}}$ and $L_{2}$ has a $\hat{0}_{L_{2}}$. By lem. 2.7, both $r_{1}$ and $r_{2}$ are positive. Thus $L_{1}$ and $L_{2}$ are positive. ∎
Remark 2.10.
(1) A McLarnan map (def. 3.1(C)) on $L$ may not factor into McLarnan maps on $L_{1}$ and $L_{2}$ in any nice way.
(2) The trivial lattice is a Fomin lattice: Its $w$ must be the empty function and $r=0$. It is the identity of the cartesian product of Fomin lattices. By our definition, this is not a positive weighting. Thus lem. 2.9(2) requires the hypothesis that $L_{1}$ and $L_{2}$ are not the trivial lattice. [^6] Though our presentation of the theory might be made more uniform if we defined the trivial Fomin lattice to be positive.
(3) Lem. 2.9(2) requires the hypothesis that $\hat{0}_{L}$ exists: Consider $\mathbb{Y}$, the lattice of partitions, which is a positive Fomin lattice with $w=1$ and $r=1$. Consider $\mathbb{Y}^{*}$, the lattice dual of $\mathbb{Y}$, which is a Fomin lattice with $w=1$ and r=1r=-1 and so is not positive. Define $2\mathbb{Y}$ to be the Fomin lattice $\mathbb{Y}$ with $w=2$ and $r=2$, which is positive. Then $L=2\mathbb{Y}\times\mathbb{Y}^{*}$ is a Fomin lattice with $w=1\textup{ or }2$ and $r=1$, which is positive. Thus $L=2\mathbb{Y}\times\mathbb{Y}^{*}$ is a positive Fomin lattice that is a product of Fomin lattices, one of which is not positive.
Cartesian product of distributive Fomin lattices
The results in this paragraph are only relevant under the assumption that $L$ is distributive and that its poset of join-irreducibles is unique-cover-modular. Hence, they are not significant until para. 5. However, since these results are not new, we place them in this section.
Lemma 2.11.
Let $L$ be a finitary777“Finitary” means that all principal ideals are finite. For lattices, it is equivalent to $\hat{0}$ exists and the lattice is locally finite. distributive lattice with poset of join-irreducible elements $P=J(L)$. Then $L$ is the cartesian product of two non-trivial lattices $L=L_{1}\times L_{2}$ iff $P$ is the disjoint union of two non-empty posets $P=P_{1}\sqcup P_{2}$, in which case $L_{1}={\operatorname{Ideal}_{\textup{f}}}(P_{1})$ and $L_{2}={\operatorname{Ideal}_{\textup{f}}}(P_{2})$. [^8] We define ${\operatorname{Ideal}_{\textup{f}}}(P)$, where $P$ is a poset, to be the lattice of finite lower order ideals of $P$.
Lemma 2.12.
Let $P$ be a poset and $P=P_{1}\sqcup P_{2}$. Then
  1. (1)
    If $P$ is cover-modular, then $P_{1}$ and $P_{2}$ are cover-modular. (def. 2.1.)
  2. (2)
    If $P$ is unique-cover-modular, then $P_{1}$ and $P_{2}$ are unique-cover-modular. (def. $2.$)
Remark 2.13.
Suppose $L$ is a distributive lattice, $L$ is finitary, and $L$ has finite upward covers. It is not necessary that $P$ has finite upward covers, as is shown by the counterexample shown in fig. 5.
(a) Lattice $L$ has finite upward covers
(b) Poset $P=J(L)$ does not have finite upward covers
Figure 5. Counterexample in rem. 2.13
Define $P$ to be the poset with elements $x,y_{0},y_{1},y_{2},\ldots,w_{0},w_{1},w_{2},\ldots$ and the coverings:
  • $w_{0}\lessdot w_{1}\lessdot w_{2}\lessdot\cdots$,
  • $x\lessdot y_{i}$ for all $i\geq 0$, and
  • $w_{i}\lessdot y_{i}$ for all $i\geq 0$.
Define $L$ to be the lattice $L={\operatorname{Ideal}_{\textup{f}}}(P)$ of all finite ideals of $P$. By Birkhoff’s Representation Theorem[Birk1967a]*§III.3 Th. 3[Stan2012a]*Prop. 3.4.3 $L$ is distributive, $L$ is finitary, and $P$ is poset-isomorphic to the set of join-irreducible elements of $L$. The elements of $L$ are (where $a,b,c,\ldots]$ is the ideal of $P$ generated by the elements $a,b,c,\ldots$):
  • $w_{i}]=\{w_{0},w_{1},w_{2},\ldots,w_{i}\}$ for all $i\geq-1$ (generating $\emptyset$ when i=1i=-1),
  • $x,w_{i}]=\{x,w_{0},w_{1},w_{2},\ldots,w_{i}\}$ for all $i\geq-1$ (generating $\{x\}$ when i=1i=-1), and
  • $w_{i},y_{n_{1}},y_{n_{2}},y_{n_{3}},\ldots y_{n_{j}}]=\{x,w_{0},w_{1},w_{2},\ldots,w_{i},y_{n_{1}},y_{n_{2}},y_{n_{3}},\ldots y_{n_{j}}\}$ for all $i\geq 0$, all $j\geq 1$, and all $0\leq n_{1}<n_{2}<n_{3}<\cdots<n_{j}\leq i$ (some non-empty finite set of $y_{\bullet}$ added to a compatible ideal of the preceding type),
We see that each of the above types of elements of $L$ (ideals of $P$) has only a finite number of ideals that cover it in $L$, that is, contain one more element than it does. So $L$ has finite upward covers. We see that $x$ has all of the $y_{\bullet}$ as upward covers in $P$. So $P$ does not have finite upward covers.
3. Fomin lattice essentials
The celebrated Robinson–Schensted algorithm[Rob1938a][Schen1961a][Fom1994a][Fom1995a] and its variants that have attracted substantial attention can be constructed[Wor2023c]*sec. 2 from the following machinery. [^9] Though since a Robinson–Schensted variant can be constructed from any DGG, it is not clear why DGGs that are not Fomin lattices have not been as productive. See item (10) in sec. 8. The core of the machinery is a “Fomin lattice” which satisfies properties (A)(G) of para. 2, summarized here:
Definition 3.1.
  1. (A)
    A Fomin lattice, [^10] We call it a Fomin lattice because all of the required properties are explicitly or implicitly given in [Fom1994a] and [Fom1995a]. which comprises:
    1. (1)
      a lattice $L$ that
      1. (a)
        is modular,
      2. (b)
        is locally finite, and
      3. (c)
        has finite upward and downward covers for each element;
    2. (2)
      a differential degree $r$ which is a member of a set of values $V$;
    3. (3)
      a weighting $w(\bullet\lessdot\bullet)$, which is a function from coverings (or prime quotients), pairs $x\lessdot y$ of elements of $L$, to $V$, for which $L$ is a weighted-differential lattice[Stan1990a]*sec. 3[Fom1994a]*sec 2.2:
      $\sum_{y\,|\,y\lessdot x}w(y\lessdot x)+r=\sum_{z\,|\,x\lessdot z}w(x\lessdot z)\textup{\qquad\qquad for all }x\in L$ (7)
      (which is often described as $r$-differential); and
    4. (4)
      for which $w$ is projective-constant, that is, constant on projectivity classes[Birk1967a]*§I.7 Def. of coverings, or equivalently, on coverings which are cover-projective: if $x\vee y$ covers $x$ and $y$ and $x\wedge y$ is covered by $x$ and $y$, then $w(x\wedge y\lessdot x)=w(y\lessdot x\vee y)$ and $w(x\wedge y\lessdot y)=w(x\lessdot x\vee y)$.
  2. (B)
    The lattice $L$ has a minimum element, $\hat{0}$.
  3. (C)
    A McLarnan map [^11] We name the correspondence for McLarnan, who seems to have been the first to emphasize[McLar1986a]*Sec. 1 p. 22 the arbitrariness of the choice of the correspondence and thus that there is no “natural” correspondence. The set of possible correspondences for a particular Fomin lattice is usually uncountable. We choose “map” because of its consonance with “McLarnan”. or $R$-correspondence $M$, which is a realization of the weighted-differential condition:
[…]
  4. (D)
    Depending on the degree of abstraction of the discussion, the requirements for the set of values $V$ varies. The minimum requirement is that $V$ be a module over a unitial ring of scalars $S$. Usually we will be unconcerned with the details of $S$ other than it necessarily contains a homomorphic image of $\mathbb{Z}$ and thus we can multiply elements of $V$ by integer coefficients. If $V=\mathbb{Z}$ (which implies $\mathbb{Z}$ is isomorphically embedded in $S$), $r>0$, and all of the values of $w$ are positive, then the weighting is defined to be positive, which is required to construct a Robinson–Schensted algorithm.
Lemma 3.2.
If a lattice $L$ forms a Fomin lattice with two pairs of differential degrees and weightings $(r,w)$ and $(r^{\prime},w^{\prime})$, then $L$ forms a Fomin lattice with any linear combination of the two: the differential degree $r^{\prime\prime}=\alpha r+\beta r^{\prime}$ and weighting $w^{\prime\prime}(x\lessdot y)=\alpha w(x\lessdot y)+\beta w^{\prime}(x\lessdot y)$ for any $\alpha,\beta\in S$. Thus, the set of pairs of degrees and weightings for $L$ forms a subspace of an infinite-dimensional module over $S$, the subspace of solutions of the set of linear equations (7).
If $(r,w)$ and $(r^{\prime},w^{\prime})$ are positive, $\alpha\geq 0$, $\beta\geq 0$, and not both $\alpha$ and $\beta$ are 0, then $(r^{\prime\prime},w^{\prime\prime})$ is positive.
Lemma 3.3.
Because $L$ is modular and $w$ satisfies the projectivity condition $(A)$, the relevant version of the Jordan–Hölder theorem[Birk1967a]*§III.7 Th. 9 Cor.[Nat2017a]*Th. 9.7 shows that $L$ is graded and that the multiset of the weights of the coverings composing a path from some $x$ to some $y\geq x$ in $L$ depends only on $x$ and $y$ and not on the particular path.
Definition 3.4.
We define the number of colorings of $x\in L$, denoted by $c(x)$, to be the product of the weights of the coverings composing an arbitrary path from $\hat{0}$ to $x$ in $L$.
Theorem 3.5.
Given this machinery, the general Robinson–Schensted insertion algorithm can be constructed using Fomin’s growth diagram construction provided the weighting is positive. The set of permutations of $\{1,\ldots,n\}$ whose elements are $r$-colored can be bijectively mapped to the set
$\{(P,Q,c)\,|\,$ there exists $x\in L$ of grade $n$ for which
[…]
where the particular bijection depends on the McLarnan map that is chosen. [^12] We use “path” rather than the more usual “saturated chain” to emphasize the importance of the coverings composed of adjacent elements of the chain, which bear the weights.
Definition 3.6.
Each element $x$ of $L$ with grade $n$ is called a diagram of size $n$. Each path from $\hat{0}$ to $x$ is called a tableau of shape $x$ and thus size $n$. We define $f(x)$ as the number of paths from $\hat{0}$ to $x$, that is, the number of tableaux of shape $x$.
Theorem 3.7.
The most celebrated consequence of the general insertion algorithm is the immediate consequence of the above bijection:
$n!\,r^{n}=\sum_{\textup{diagrams $x$ with size $n$}}f(x)^{2}c(x).$ (8)
4. Auxiliary result
The result here supports the discussion of future directions (sec. 8). It is placed here to avoid interacting with the assumptions used in sec. 5 and 6.
Partitioning a lattice
Let $L$ be a locally finite modular lattice and let $S$ be a set of coverings in $L$ that is closed under cover-projectivity, that is, if $x\lessdot y$ is in $S$ and it is cover-projective to $x^{\prime}\lessdot y^{\prime}$, then $x^{\prime}\lessdot y^{\prime}$ is in $S$. Since $L$ is locally finite and modular, this is equivalent to that $S$ is the union of a set of projectivity classes of coverings in $L$.
We will define a relationship $\sim$ between elements of $L$. Intuitively, $x\sim y$ if there is a (finite) path between $x$ and $y$ composed of (upward and downward) covering steps, none of which are in $S$. But it’s easier to analyze $\sim$ by defining it in a different way.
Lemma 4.1.
Given $x\leq y$ in $L$ and any two paths between them, one path contains a covering in $S$ iff the other path contains a covering in $S$.
Proof.
Since $L$ is modular and locally finite, the interval $[x,y]$ is modular and finite. Choose any projectivity class $C$ of coverings in $L$. The relevant version of the Jordan–Hölder theorem[Birk1967a]*§III.7 Th. 9 Cor.[Nat2017a]*Th. 9.7 can be applied $[x,y]$ to show that one path contains a covering in $C$ iff the other path contains a covering in $C$. The lemma then follows. ∎
Thus the following definition does not depend on the particular path chosen between $x\wedge y$ and $x\vee y$.
Definition 4.2.
We define $x\sim_{S}y$ (or $x\sim y$ if, as usual, $S$ is understood) iff no covering in a path from $x\wedge y$ to $x\vee y$ is in $S$.
Lemma 4.3.
If $x\leq y$ in $L$, then $x\sim y$ iff no covering in a path from $x$ to $y$ is in $S$.
Lemma 4.4.
If $x\leq y\leq z$ in $L$ and $x\sim z$, then $x\sim y$ and $y\sim z$.
Proof.
By lem. 4.3, no covering in any path from $x$ to $z$ is in $S$. We can choose such a path from $x$ to $y$ and from there to $z$. Since no covering in that path is in $S$, no covering from $x$ to $y$ or from $y$ to $z$ is in $S$. Then by lem. 4.3, $x\sim y$ and $y\sim z$. ∎
Lemma 4.5.
Conversely, if $x\leq y\leq z$ in $L$, $x\sim y$, and $y\sim z$, then $x\sim y$.
Lemma 4.6.
If $x,y\in L$, then $x\sim x\vee y$ iff $x\wedge y\sim y$.
Proof.
This follows from the diamond isomorphism theorem and lem. 4.3. ∎
Lemma 4.7.
If $\in L, xyx\sim y$ implies $x\wedge y\sim x\vee y$, $x\wedge y\sim x\sim x\vee y$, and $x\wedge y\sim y\sim x\vee y$.
Lemma 4.8.
If $\in L, xxyy$$\sim x\vee$ $\sim y$ implies $x\sim y$. Dually, $x\sim x\wedge y\sim y$ implies $x\sim y$.
Proof.
Given $x\sim x\vee y$, by lem. 4.6, $x\wedge y\sim y$. Since $x\wedge y\leq y\leq x\vee y$, $x\wedge y\sim y$, and $y\sim x\vee y$, By lem. 4.5, $x\wedge y\sim x\vee y$ and so $x\sim y$. The dual is proved dually. ∎
Lemma 4.9.
Given $x,y,z\in L$, if $x\sim y$ and $y\sim z$, then $x\sim z$.
Proof.
The construction for this proof is shown in fig. 6.
Figure 6. Construction for the proof of lem. 4.9.
We know $x\sim y$ and $y\sim z$. By lem. 4.7, $x\sim x\vee y$, $x\vee y\sim y$, $y\sim y\vee z$, and $y\vee z\sim z$. Define $v=(x\vee y)\wedge(y\vee z)$. Of necessity, $y\leq v\leq x\vee y$ and $y\leq v\leq y\vee z$. By lem. 4.4, $v\sim x\vee y$ and $v\sim y\vee z$. By lem. 4.8, $x\vee y\sim y\vee z$ and by lem. 4.7, $x\vee y\sim(x\vee y)\vee(y\vee z)=x\vee y\vee z$. Since $x\leq x\vee y\leq x\vee y\vee z$, by lem. 4.5, $x\sim x\vee y\vee z$. Since $x\leq x\vee z\leq x\vee y\vee z$, by lem. 4.4, $x\sim x\vee z$. Similarly, $z\sim x\vee z$. Then by lem. 4.8, $x\sim z$. ∎
Lemma 4.10.
$\sim$ is an equivalence relation on the elements of $L$.
Proof.
That $\sim$ is reflexive and symmetric are trivial. Lem. 4.9 shows that $\sim$ is transitive. ∎
Lemma 4.11.
Given $x,y,z\in L$, if $x\sim y$ and $x\sim z$, then $x\sim y\wedge z$ and $x\sim y\vee z$.
Proof.
By lem. 4.9, $y\sim z$, and by lem. 4.7, $y\sim y\wedge z$. Then by lem. 4.9, $x\sim y\wedge z$. Dually, $x\sim y\vee z$. ∎
Lemma 4.12.
Given $x,y,z,v\in L$ with $y\leq z\leq v$, if $x\sim y$ and $x\sim v$, then $x\sim z$.
Proof.
By lem. 4.9, $y\sim v$, and by lem. 4.4, $y\sim z$. Then by lem. 4.9, $x\sim z$. ∎
Now we fix an element $\in L. zz$ will usually be $\hat{0}$.
Definition 4.13.
We define $L_{/Sz}$ to be the equivalence class of $\sim_{S}$ containing $z$, which is $\{x\in L\,|\,x\sim_{S}z\}$.
Intuitively, we construct $L_{/Sz}$ by constructing the Hasse diagram of $L$, removing from it the edges which are coverings in $S$, selecting from the resulting graph the connected component that contains $z$, and considering that component as the Hasse diagram of a lattice, which is $L_{/Sz}$.
Lemma 4.14.
$L_{/Sz}$ is a convex sublattice of $L$.
Proof.
That $L_{/Sz}$ is convex is shown by lem. 4.12. That $L_{/Sz}$ is a sublattice is shown by lem. 4.11. ∎
We now apply this construction to a Fomin lattice that has a non-positive weighting.
Definition 4.15.
Given a Fomin lattice $L$ with weighting $w$ and $\hat{0}$, we define its truncation $L_{/}$ to be $L_{/S\hat{0}}$, where $S=\{x\lessdot y\,|\,w(x\lessdot y)=0\}$, the set of coverings with weight 0. We define the weighting $w_{/}$ on $L_{/}$ to be $w$ restricted to $L_{/}$.
Theorem 4.16.
Given a Fomin lattice $L$ with $\hat{0}$, its truncation $L_{/}$ is a Fomin lattice with (the same) $\hat{0}$.
Proof.
All of the requirements for $L_{/}$ to be a Fomin lattice (def. 3.1) are trivially inherited from $L$ except for (A). But by construction, equation (7) for a particular $x\in L_{/}$ contains a subset of the terms of (7) for $x\in L$, and the terms of the latter that are not in the former are ones for which $w$ has the value 0. Thus (7) is true for all $x\in L_{/}$. ∎
Theorem 4.17.
Given a finitary distributive Fomin lattice $L$ with $\hat{0}$ with points $J(L)$, its truncation $L_{/}$ is isomorphic to ${\operatorname{Ideal}_{\textup{f}}}(P_{/})$ where $P_{/}$ is the subposet $\{j\in J(L)\,|\,\textup{not }(\exists k)(k\leq j\textup{ and }w(j)=0)\}$.
Theorem 4.18.
Given a Fomin lattice $L$ with $\hat{0}$, if its set of values is $\mathbb{Z}$, $r>0$, and for every covering $x\lessdot y$ for which $w(x\lessdot y)<0$, there exists a $z,v$ for which $z\lessdot v\leq x$ and $w(z\lessdot v)=0$, then $L_{/}$ is a Fomin lattice with $\hat{0}$ and a positive weighting.
Proof.
By th. 4.16, $L_{/}$ is a Fomin lattice with $\hat{0}$. It remains to prove that its weighting $w_{/}$ is has only positive values on $L_{/}$. Assume to the contrary that $x\lessdot y$ in $L_{/}$ and $w_{/}(x\lessdot y)\leq 0$. If $w_{/}(x\lessdot y)=0$, then by the construction of $L_{/}$ and lem. 4.3, $y\not\in L_{/}$. If on the other hand $w_{/}(x\lessdot y)<0$, then by hypothesis there is $\hat{0}\leq z\lessdot v\leq y$ with $w_{/}(z\lessdot v)=0$, and similarly $v\not\in L_{/}$, implying $y\not\in Z_{/}$. ∎
Theorem 4.19.
Given a distributive Fomin lattice $L$ with $\hat{0}$ with points $P(L)$, if for every point $j$ for which $w(j)<0$ then there exists a point $k$ for which $k\leq j$ and $w(k)=0$, then $L_{/}$ is a Fomin lattice with a positive weighting.
Of course, this construction of a smaller lattice that is an positive Fomin lattice could be done by choosing the component containing a different element $z\in L$ than $\hat{0}$, since if $L$ has zerohatzerohat, then any sublattice of $L$ has a minimum element.
The only known application of truncation is when $L$ is $\mathbb{Y}$, $P=J(L)$ is the quadrant, and the weighting on $P$ is $w((i,j)=\alpha+\beta i+\gamma j$.[Fom1994a]*Exam. 2.2.7 If $\alpha=r$, $\beta=r/k$, and $\gamma=-r/k$ for some nonzero integer $k$ that divides $r$, then $L_{/}$ is $\mathbb{Y}_{k}$ with the weights multiplied by $r/k$ (if $k$ is positive) or its left-right reversal (if $k$ is negative).
5. Local characterization
We start with proving various “local” facts about the poset $P$ of points (join-irreducible elements) of $L$. These culminate with a classification of the possible structures of the neighborhood of a point $p$, that is, the convex set generated by the points that cover of $p$ and the points that are covered by $p$.
Henceforth we will characterize positive Fomin lattices under certain sets of assumptions. As we proceed through the analysis, further assumptions will be added. All of these assumptions may seem an excessively narrow approach, but all known Fomin lattices with $\hat{0}$ other than $\mathbb{YF}_{r}$ satisfy all of these assumptions.
Distributive lattice
Assumption 5.1.
Henceforth, we will restrict our attention to distributive Fomin lattices with $\hat{0}$. Specifically, we assume $L$ is a Fomin lattice with $\hat{0}$ with differential degree $r$, values $V$, and weighting $w$, and that $L$ is distributive. (We do not, at this point, assume that $w$ is positive.)
Theorem 5.2.
[Birk1967a]*§III.3 Th. 3[Stan2012a]*Prop. 3.4.3 Since we have assumed that $\hat{0}$ exists, $L$ is finitary, that is, all if its principal ideals are finite. $L$ is isomorphic to the set of finite order ideals of the poset $P=J(L)$ of its join-irreducible elements, which we denote $L={\operatorname{Ideal}_{\textup{f}}}(P)$. Given a covering $x\lessdot y$ in $L$, $x\setminus y$ (considered as the set difference of two ideals of $P$) contains exactly one element of $P$, and the projectivity requirement (condition $(A)$) on $w$ can be simplified to: There is a weighting function $w$ from $P$ into $V$, and for every $x\lessdot y$ in $L$, $w(x\lessdot y)=w(y\setminus x)$.
The weighted-differential condition (7) is equivalent to
[…]
Because the image of $w$ in $L$ is the same as the image of $w$ on $P$, we abuse language by conflating the two meanings of $w$ when we are dealing them as entire functions. In particular, each of them has all of its values positive integers iff the other does, and so “$w$ is positive” is used to mean both.
Definition 5.3.
The elements of $P$ are called points or boxes.
Lemma 5.4.
  1. (1)
    $P$ is locally finite.
  2. (2)
    $P$ is finitary.
Proof.
Regarding (1): If there was some $a\leq b$ in $P$ for which $[a,b]$ is infinite, then the set of all (closed) principal ideals of the elements in $[a,b]$, $\{[x]\,|\,a\leq x\leq b\}$, would be an infinite subset of $L$ with all of its elements $\geq[a]$ and $\leq[b]$. That contradicts that $L$ is locally finite.
Regarding (2): Similarly, if there was an $a\in P$ whose principal ideal $[a]$ in $P$ was infinite, then $\{[x]\,|\,x\leq b\}$ would be an infinite subset of $L$ with all of its elements $\leq[a]$. That contradicts that, given our assumptions, th. 5.2 shows that $L$ is finitary. ∎
Lemma 5.5.
The lattice of diagrams is isomorphic to the lattice of finite ideals of points under union and intersection, the size of a diagram is the cardinality of its corresponding ideal (the number of points it contains), and the number of tableaux with a particular shape is the number of linear orderings of its points (a subposet of $P$). The number of colorings $c(x)$ of a diagram $x$ is the product of the weights $w$ of its points: $c(x)=\prod_{a\in x}w(a)$.
Definition 5.6.
Given any diagram $x$, its insertion points are the points that are the minimal elements of $P\setminus x$, those which can be added to $x$ to make a diagram (whose size is 1 larger than the size of $x$). The deletion points of $x$ are the points that are the maximal elements of $x$, those which can be deleted from $x$ to make a diagram (whose size is 1 smaller than the size of $x$). For any diagram $x$, we define $D_{x}$ to be the set of deletion points of $x$, and $I_{x}$ to be the set of insertion points of $x$.
The summation on the left of (9) is over the deletion points of $x$ and the summation on the right of (9) is over the insertion points of $x$. Thus, (9) is equivalent to
$\sum_{a\in D_{x}}w(a)+r=\sum_{b\in I_{x}}w(b)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$}.$ (10)
Applying (9) to the ideal $\emptyset$ gives
$r=\sum_{z\,|\,z\textup{ minimal in }P}w(z)$ (11)
In many cases $P$ has a unique minimal element, $\hat{0}_{P}$, and in that case, (11) reduces to $w(\hat{0}_{P})=r$.
Given an ideal $x$ and a point $p$ in $I_{x}$, that is, minimal in $P\setminus x$, we can construct the ideal ${x^{\prime}}=x\sqcup\{p\}$. Thus $P\setminus x=(P\setminus x^{\prime})\sqcup\{p\}$ and $p\in D(x^{\prime})$.
The difference between (10) applied to $x^{\prime}$ and (10) applied to $x$ is
$\sum_{a\in D_{x^{\prime}}\setminus D_{x}}w(a)-\sum_{b\in D_{x}\setminus D_{x^{\prime}}}w(b)=\sum_{a\in I_{x^{\prime}}\setminus I_{x}}w(a)-\sum_{b\in I_{x}\setminus I_{x^{\prime}}}w(b)\textup{\qquad\qquad for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$, $p\in I_{x}$}.$ (12)
The domains of the four summations of (12) are:
  1. (1)
    $D_{x^{\prime}}\setminus D_{x}$: The set of maximal elements of $x^{\prime}$} that are not maximal elements of $x$, which is $\{p\}$,
  2. (2)
    $D_{x}\setminus D_{x^{\prime}}$: The set of maximal elements of $x$ that are not maximal elements of $x^{\prime}$},
  3. (3)
    $I_{x^{\prime}}\setminus I_{x}$: The set of minimal elements of $P\setminus{x^{\prime}}$ that are not minimal elements of $P\setminus x$,
  4. (4)
    $I_{x}\setminus I_{x^{\prime}}$: The set of minimal elements of $P\setminus x$ that are not minimal elements of $P\setminus{x^{\prime}}$, which is $\{p\}$.
Lemma 5.7.
Condition (9) (for every $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$) is equivalent to the combination of (11) and (12) (for every $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$ and $p\in I_{x}$).
Proof.
Regarding $\Rightarrow$: This has been proved by the above derivation of (11) and (12) from (9).
Regarding $\Leftarrow$: For any $y\in{\operatorname{Ideal}_{\textup{f}}}(P)$, we want to prove (9) for its $x=y$. Since $y$ is finite, we can prove this by induction upward on the size of $y$. The base case is $y=\emptyset=\hat{0}$. In that case, (9) with its $x=\hat{0}$ reduces to (11).
If $y$ has positive size, let $y^{\prime}\lessdot y$ in ${\operatorname{Ideal}_{\textup{f}}}(P)$. By induction, (9) is true for its $x=y^{\prime}$. Now consider (12) with its $x=y^{\prime}$ and its $p$ the unique element of $y\setminus y^{\prime}$. We can add these two equations and by the above derivation obtain (9) for its $x=y$. ∎
Definition 5.8.
Given a point $p\in P$, define $C^{-}_{p}$ ($C^{-}$ for “downward covers”) as the set of all elements of $P$ which are covered by $p$. Similarly, define $C^{+}_{p}$ ($C^{+}$ for “upward covers”) as the set of all elements of $P$ which cover $p$. Define $N_{p}$ (NN for “neighborhood”) to be $\{x\in P\,|\,(\exists a\in C_{p}^{-},b\in C_{p}^{+})\,a\leq x\leq b\}$, the convex subset of $P$ generated by $C_{p}^{-}$ and $C_{p}^{+}$. Define $S_{p}$ ($S$ for “siblings”) to be $\{x\in P\,|\,x\neq p\textup{ and }(\exists a\in C_{p}^{-},b\in C_{p}^{+})\,a<x<b\}=N_{p}\setminus C_{p}^{-}\setminus C_{p}^{+}\setminus\{p\}$, $N_{p}$ with $C_{p}^{-}$, $C_{p}^{+}$, and $p$ deleted.
Lemma 5.9.
By construction, every element of $C_{p}^{+}$ is >> every element of $C_{p}^{-}$, but no element of $C_{p}^{+}$ covers any element of $C_{p}^{-}$. All elements of $C_{p}^{+}$ are incomparable to each other. Dually for $C_{p}^{-}$. Every element of $S_{p}$ is incomparable to $p$. If $P$ is graded, every element of $S_{p}$ has the same grade as $p$ and $S_{p}$ is an antichain. However, in general, $S_{p}$ need not be an antichain.
Lemma 5.10.
Given an ideal $x$ and a point $p$ that is an insertion point of $x$, $C_{p}^{-}\subset x$ and $C_{p}^{+}\subset P\setminus(x\sqcup\{p\})$.
Definition 5.11.
Given an ideal $x$ and a point $p$ in $I_{x}$, define $S^{-}_{px}=S_{p}\cap x$ and $S^{+}_{px}=S_{p}\cap(P\setminus(x\sqcup\{p\}))$.
Lemma 5.12.
Since $p\not\in S$, by construction, $S_{p}=S_{px}^{-}\sqcup S_{px}^{+}$, and if $s^{-}\in S_{px}^{-}$ and $s^{+}\in S_{px}^{+}$ are comparable, then $s^{-}\leq s^{+}$. So $S_{px}^{-}$ is an ideal of $S_{p}$ and $S_{px}^{+}$ is a filter of $S_{p}$.
The generic relationships of the elements of these sets are shown in fig. 7.
Figure 7. Adding a point $p$ to an ideal $x$ making an ideal $x^{\prime}$
Unique-cover-modular
Assumption 5.13.
Henceforth, we assume the poset of points is unique-cover-modular (def. 2.1).
The following example disproves a conjecture that was in preliminary versions of this article.
Remark 5.14.
The poset $t$-Plait in [Fom1994a]*Exam. 2.2.10 and Fig. 7 is locally finite and cover-modular but is not graded.131313“Graded” means that every saturated chain from $a$ to $b$ has the same length, but there need not be a $\hat{0}$. In general, the grade function of a graded poset ranges over the integers, not just the nonnegative integers.
Lemma 5.15.
Let $p\in P$. For any $a\in C_{p}^{+}$ and any $b\neq p$ that is covered by $a$, unique-cover-modularity shows that there is a $c\in C_{p}^{-}$ that is covered by $a$ and so $b\in S_{p}$. Dually, for any $a\in C_{p}^{-}$ and any $b\neq p$ that covers $a$, there is a $c\in C_{p}^{+}$ that covers $a$ and so $b\in S_{p}$. For any $a\in S_{p}$, unique-cover-modularity shows there is a unique element of $C_{p}^{+}$ which is >a>a. Dually, there is a unique element of $C_{p}^{-}$ which is <a<a.
Definition 5.16.
For any $p\in P$ and any $a\in S_{p}$, we define $a^{+}_{p}$ to be the unique element of $C_{p}^{+}$ which is >a>a. Dually, we define $a^{-}_{p}$ to be the unique element of $C_{p}^{-}$ which is <a<a.
Lemma 5.17.
The four summation domain sets of (12) are:
  1. (1)
    $D_{x^{\prime}}\setminus D_{x}=\{p\}$
  2. (2)
    $D_{x}\setminus D_{x^{\prime}}$ is the set of maximal elements of $x$ which are covered by $p$. This is the same as the elements of $C_{p}^{-}$ which are not $\leq$ (or equivalently, not covered by) any element of $S_{px}^{-}$. We denote this property by $\not\leq S_{px}^{-}$.
  3. (3)
    $I_{x^{\prime}}\setminus I_{x}$ is the set of minimal elements of $P\setminus x^{\prime}$ which cover $p$. This is the same as the elements of $C_{p}^{+}$ which are not $\geq$ (or equivalently, do not cover) any element of $S_{px}^{+}$. We denote this property by $\not\geq S_{px}^{+}$.
  4. (4)
    $I_{x}\setminus I_{x^{\prime}}=\{p\}$
This shows (12) is equivalent to
$w(p)-\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq S_{px}^{-}}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq S_{px}^{+}}w(b)-w(p)$ for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$, $p\in I_{x}$
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq S^{-}_{px}}w(a)+\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq S^{+}_{px}}w(b)$ $=2w(p)$ $\textup{for all $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$, $p\in I_{x}$}.$ (13)
Proof.
Regarding (1) and (4): These are straightforward.
Regarding (2) $\Rightarrow$: Let $a\in D_{x}\setminus D_{x^{\prime}}$. Then $a$ is maximal in $x$ but not maximal in $x^{\prime}$}. Thus $a\leq$ an element of $x^{\prime}$ which is not an element of $x$. The only element of $x^{\prime}$ that is not an element of $s$ is $p$, so $a<p$. If $a$ is not covered by $p$, then there is a $b$ with $a<b<p$. Thus $b\in x^{\prime}$ but since $a$ is maximal in $x$, $b\not\in x$, which requires that $b=p$, which is a contradiction. Thus $a$ is covered by $p$ and so $a\in C^{-}_{p}$.
To show that $a\not\leq S_{px}^{-}$, assume there is a $c$ in $S_{px}^{-}$ with $a\leq c$. Since $C^{-}_{p}$ is disjoint from $S^{-}_{ps}$, $a<c$. But $c\in x$, contradicting that $a$ is maximal in $x$. Thus $a\not\leq S_{px}^{-}$.
Regarding (2) $\Leftarrow$: Let $a\in C^{-}_{p}$ and $a\not\leq S^{-}_{px}$. Because $a<p$ and $p\in x^{\prime}$, $a$ is not maximal in $x^{\prime}$.
To show that $a$ is maximal in $x$, assume there is a $d$ in $x$ with $d>a$. We can assume that $d\gtrdot a$. Then we apply unique-cover-modularity to $p$ and $d$ to obtain a $e$ which covers $p$ and $d$. So $e\in C^{+}_{p}$, and with $a\in C^{-}_{p}$ shows $d\in S_{p}$. Since $d\in x$, $d\in S^{-}_{px}$. This contradicts that $a\not\leq S^{-}_{ps}$, so $a$ is maximal in $x$.
Regarding (3): This is proved dually to (2). ∎
Reorganizing the linear equations
Looking at (13) as a set of linear equations in the values of $w$, we see that each equation is determined by choosing first an $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$ and then a $p\in I_{x}$. We will now reorganize the statement of the set of linear equations (13), but indexed first by an arbitrary choice of $p\in P$ (which determines $C_{p}^{+}$ and $C_{p}^{-}$), and then by the values of $S_{px}^{+}$ and $S_{px}^{-}$ consistent with $p$.
Lemma 5.18.
The set of equations (13) is the same set as the equations
[…]
Proof.
A particular equation in the set (13) is indexed first by an $x\in{\operatorname{Ideal}_{\textup{f}}}(P)$ and then a $\in I_{x}. pp$ determines $C_{p}^{+}$ and $C_{p}^{-}$, and $p$ and $x$ together determine $S^{-}_{px}$ and $S^{+}_{px}$. Necessarily, $S_{p}=S^{+}_{px}\sqcup S^{+}_{px}$, so setting $T^{-}=S^{-}_{px}$ and $T^{+}=S^{+}_{px}$ shows that the equation is in the set (14).
Conversely, a particular equation in the set (14) is indexed first by a $p\in P$ and then a partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$. Define $x=\{z\in P\,|\,z<p\}\sqcup T^{-}=(p)\sqcup T^{-}$, the principal ideal of $P$ generated by $p$, less $p$ itself, plus $T^{-}$. By construction, $x$ is an ideal of $P$, $p$ is in insertion point of $x$, $S^{-}_{px}=T^{-}$ and $S^{+}_{px}=T^{+}$. This shows that the equation is in the set (13). ∎
Orphans
Definition 5.19.
Define $x\in C_{p}^{+}$ as an upward orphan (relative to $p$) if $x$ does not cover any element of $S_{p}$, which is equivalent to $x$ covering only one element of $N_{p}$ (namely $p$). Similarly, we define $x\in C_{p}^{-}$ as a downward orphan (relative to $p$) if $x$ is not covered by any element of $S_{p}$, which is equivalent to $x$ being covered by only one element of $N_{p}$ (namely $p$).
In (14), if we choose $T^{-}=\emptyset$ and $T^{+}=S_{p}$, it becomes
$\sum_{a\in C_{p}^{-}}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}}w(b)$ $=2w(p)\textup{\qquad for all $p\in P$}$
By lem. 5.15, $S_{p}$ contains every element covered by any element of $C_{p}^{+}$ except for $p$, so the second sum is over all upward orphans. Thus
Lemma 5.20.
$\sum_{a\,|\,a\lessdot p}w(a)+\sum_{b\,|\,b\gtrdot p,\ b\textup{ upward orphan}}w(b)$ $=2w(p)\textup{\qquad for all $p\in P$}$ (15)
Equation (15) allows the weight of a point to be computed as half of the sum of the weights of the elements it covers, if the point is not covered by any upward orphans.
If we choose $T^{-}=S_{p}$ and $T^{+}=\emptyset$, then (14) becomes
Lemma 5.21.
$\sum_{a\,|\,a\lessdot p,\ a\textup{ downward orphan}}w(a)+\sum_{b\,|\,b\gtrdot p}w(b)$ $=2w(p)\textup{\qquad for all $p\in P$}$ (16)
Partitioning $S$ into three sets
Given $p\in P$, let us partition $S_{p}$ into three disjoint parts $S_{p}=L\sqcup\{m\}\sqcup U$ where $L$ is an ideal of $S$, $U$ is a filter of $S$, and thus $m$ is an element between $L$ and $U$. In (14), setting its $T^{-}=L$ and its $T^{+}=\{m\}\sqcup U$ gives
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L}w(a)+\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq\{m\}\sqcup U}w(b)$ $=2w(p)$ (17)
and then setting its $T^{-}=L\sqcup\{m\}$ and its $T^{+}=U$ in (14) gives
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L\sqcup\{m\}}w(a)+\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U}w(b)$ $=2w(p)$ (18)
Since $x\not\leq L\sqcup\{m\}$ implies $x\not\leq L$ and $x\not\geq\{m\}\sqcup U$ implies $x\not\geq U$, subtracting (18) from (17) gives
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L,\ \textup{not }a\not\leq L\sqcup\{m\}}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U,\ \textup{not }b\not\geq\{m\}\sqcup U}w(b)$
This is equivalent to
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L,\ a\leq m}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U,\ b\geq m}w(b)$ (19)
Lemma 5.22.
Given a $p\in P$, equation (14) holds for every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$ iff:
  1. (1)
    equation (15) holds for $p$ and
  2. (2)
    equation (19) holds for $p$ and for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$, where L is an ideal of $S_{p}$ and $U$ is a filter of $S_{p}$.
Proof.
Regarding $\Rightarrow$: This has been proved by the above derivations of (15), (16), and (19).
Regarding $\Leftarrow$: For a given $p$, we prove (14) for any partition $S_{p}=T^{-}\sqcup T^{+}$ where $T^{-}$ is an ideal and $T^{+}$ is a filter by induction upward on the size of $T^{-}$. If $T^{-}=\emptyset$, then by hypothesis (15) holds for $p$, which is equivalent to (14) for $T^{-}=\emptyset$.
If $T^{-}\neq\emptyset$, then choose $m$ as a maximal element of $T^{-}$ and define $L=T^{-}\setminus\{m\}$ and $U=T^{+}$, so that $S_{p}=L\sqcup\{m\}\sqcup U$ with $L$ being an ideal and $U$ being a filter. Since $|L|=|T^{-}|-1$, by induction (14) holds for the partition $S_{p}=L\sqcup(\{m\}\sqcup U)$, which is
$\sum_{a\in C_{p}^{-},\ a\not\leq L}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq\{m\}\sqcup U}w(b)$ $=2w(p)$ (20)
By hypothesis, (19) holds for the partition $S=L\sqcup\{m\}\sqcup U$, which is
$\sum_{a\,|\,a\in C_{p}^{-},\ a\not\leq L,\ a\leq m}w(a)$ $=\sum_{b\,|\,b\in C_{p}^{+},\ b\not\geq U,\ b\geq m}w(b)$ (21)
Subtracting the left side of (21) from the first term of (20) and adding the right side of (21) to the second term of (20) gives
$\sum_{a\in C_{p}^{-},\ a\not\leq L,\ a\not\leq m}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq U}w(b)$ $=2w(p)$
This equation is the same as (14) for the partition $S_{p}=(L\sqcup\{m\})\sqcup U=T^{-}\sqcup T^{+}$.
Thus by induction (14) holds for $p$ and every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$. ∎
We can replace (15) with (16):
Lemma 5.23.
Given a $p\in P$, equation (14) holds for every partition of $S_{p}$ into an ideal $T^{-}$ and a filter $T^{+}$ iff:
  1. (1)
    equation (16) holds for $p$ and
  2. (2)
    equation (19) holds for $p$ and for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$, where L is an ideal of $S_{p}$ and $U$ is a filter of $S_{p}$.
Proof.
This is proved dually to the proof of lem. 5.22, using (16) in place of (15). ∎
Positive weighting
Assumption 5.24.
Henceforth, we assume that $w$ is an positive weighting, that is, the set of values $V$, the target of $w$, is $\mathbb{Z}$, and all of the values of $w$ are positive.
This assumption is not a loss of generality for our larger purpose, because if a non-positive weighting could be used to construct a positive weighting using th. 4.16 et seq., our analysis would reveal the resulting positive weighting directly.
Lemma 5.25.
For any $p\in P$ and for any $x,y\in S_{p}$, if $x^{+}_{p}=y^{+}_{p}$ and $x^{-}_{p}=y^{-}_{p}$, then $x=y$.
Proof.
Assume there is $x,y\in S_{p}$, with $x\neq y$, $x^{+}_{p}=y^{+}_{p}$, and $x^{-}_{p}=y^{-}_{p}$. Since $x\neq y$, either $x\not\leq y$ or $y\not\leq x$. Without loss of generality, assume $y\not\leq x$. Define $X=\{s\in S_{p}\,|\,s^{-}_{p}=x^{-}_{p}\textup{ and }s^{+}_{p}=x^{+}_{p}\textup{ and }s\leq x\}=[x]\cap[x^{-}_{p},x^{+}_{p}$], the principal ideal generated by $x$ within the interval $x^{-}_{p},x^{+}_{p}$]. By construction, $x\in X$ and $y\not\in X$.
Given $a\in C_{p}^{-}$, $a\leq X$ iff $a\leq x$, so $a\not\leq X$ iff $a\not\leq x$. Given $b\in C_{p}^{+}$, if $b\geq$ some $z\in X$, $z^{+}_{p}=b$. But since $z\leq x$, lem. 5.15 shows $z^{+}_{p}=x^{+}_{p}$, so $b=z^{+}_{p}=x^{+}_{p}=y^{+}_{p}$ and $b\geq y\not\in X$. Thus $b\geq S_{p}$ iff $b\geq S_{p}\setminus X$ and $b\not\geq S_{p}$ iff $b\not\geq S_{p}\setminus X$.
Applying (14) to $S_{p}=\emptyset\sqcup S_{p}$ gives
$\sum_{a\in C_{p}^{-},\ a\not\leq\emptyset}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}}w(b)$ $=2w(p)$
which is equivalent to
$\sum_{a\in C_{p}^{-}}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}\setminus X}w(b)$ $=2w(p)$ (22)
Applying (14) to $S_{p}=X\sqcup(S_{p}\setminus X)$ gives
$\sum_{a\in C_{p}^{-},\ a\not\leq X}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}\setminus X}w(b)$ $=2w(p)$
which is equivalent to
$\sum_{a\in C_{p}^{-},\ a\not\leq x}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}\setminus X}w(b)$ $=2w(p)$ (23)
Subtracting (23) from (22) gives $w(x^{-}_{p})=0$. By the assumption that $w$ is positive, this is a contradiction. Thus $x=y$. ∎
Lemma 5.26.
For any $p\in P$, $S_{p}$ is an antichain.
Proof.
Assume there exists $x<y$ in $S_{p}$. By lem. 5.15, $x^{+}_{p}=y^{+}_{p}$ and $x^{-}_{p}=y^{-}_{p}$. By lem. 5.25, $x=y$, which is a contradiction. Thus $S_{p}$ is an antichain. ∎
Thus, for any $p$, any subset of $S_{p}$ is both an ideal and a filter, which allows (14) and (19) to be applied quite broadly. Consider how unique-cover-modularity simplifies (19): For any given $m\in S_{p}$, any $a$ on the left-hand side of (19) can only be $m^{-}_{p}$. Similarly, any $b$ on the right-hand side can only be $m^{+}_{p}$. Thus (19) is equivalent to:
$\begin{rcases}\textup{if }m^{-}_{p}\not\leq L\textup{ }&w(m^{-}_{p})\\ \textup{otherwise }&0\end{rcases}$ $=\begin{cases}w(m^{+}_{p})&\textup{ if }m^{+}_{p}\not\geq U\\ 0&\textup{ otherwise}\end{cases}$ (24)
If we set $L=\emptyset$ and $U=S_{p}\setminus\{m\}$,
$w(m^{-}_{p})$ $=\begin{cases}w(m^{+}_{p})&\textup{ if }m^{+}_{p}\not\geq U\\ 0&\textup{ otherwise}\end{cases}$ (25)
Given the assumption of unique-cover-modularity, $m^{+}_{p}\geq z$ for some $z\in U$ iff $m^{+}_{p}=z^{+}_{p}$ for some $z\in U$. So $m^{+}_{p}\not\geq U$ iff there is no $z\in U$ with $m^{+}_{p}=z^{+}_{p}$. Given the assumption of positivity, $w(m^{-}_{p})>0$, which is contradicted by (25) unless there is no $z\in S_{p}$ with $z^{+}_{p}=m^{+}_{p}$ and $z\neq m$.
Dually, if we set $L=S_{p}\setminus\{m\}$ and $U=\emptyset$, we get a contradiction unless there is no $z\in S_{p}$ with $z^{-}_{p}=m^{-}_{p}$ and $z\neq m$.
Thus for every $m\neq n$ in $S_{p}$, $m^{-}_{p}\neq n^{-}_{p}$ and $m^{+}_{p}\neq n^{+}_{p}$, and so the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$. In addition, (24) reduces to
$w(m^{-}_{p})$ $=w(m^{+}_{p})\textup{\qquad for all $p\in P$ and $m\in S_{p}$}$ (26)
Lemma 5.27.
Given $p\in P$, (19) holds for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$ with $L$ an ideal of $S_{P}$ and $U$ a filter of $S_{p}$ iff the following conditions are true:
  1. (1)
    the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$ and
  2. (2)
[…]
Proof.
Regarding $\Rightarrow$: This has been shown by the above derivations.
[…]
At this point, let us summarize our progress:
Lemma 5.28.
Given
  1. (assump. 5.1) $L$ satisfies the criteria of a Fomin lattice other than $(A)$, the weighted-differential property,
  2. (assump. 5.1) $L$ has a $\hat{0}$,
  3. (assump. 5.1) $L$ is distributive,
  4. (assump. 5.13) $P$ is unique-cover-modular, and
  5. (assump. 5.24) $w(\bullet)$ from $P$ into $V=\mathbb{Z}$ has only positive values,
then $L$ satisfies (9) (and thus is a Fomin lattice) iff
  1. (eq. 11) $r=\sum_{z\,|\,z\textup{ minimal in }P}w(z)$,
  2. (eq. 15) $\sum_{a\,|\,a\lessdot p}w(a)+\sum_{b\,|\,b\gtrdot p,\ b\textup{ upward orphan}}w(b)=2w(p)$ for all $p\in P$,
  3. (lem. 5.27(1)) the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$ for all $p\in P$, and
[…]
Proof.
This is proved by chaining together th. 5.2, lem. 5.7, lem. 5.17, lem. 5.18, lem. 5.22, and lem. 5.27. ∎
Similarly to lem. 5.22 and 5.23, we can replace (15) with (16):
Lemma 5.29.
Given
  1. (assump. 5.1) $L$ satisfies the criteria of a Fomin lattice other than $(A)$, the weighted-differential property,
  2. (assump. 5.1) $L$ has a $\hat{0}$,
  3. (assump. 5.1) $L$ is distributive,
  4. (assump. 5.13) $P$ is unique-cover-modular, and
  5. (assump. 5.24) $w(\bullet)$ from $P$ into $V=\mathbb{Z}$ has only positive values,
then $L$ satisfies (9) (and thus is a Fomin lattice) iff
  1. (eq. 11) $r=\sum_{z\,|\,z\textup{ minimal in }P}w(z)$,
  2. (eq. 16) $\sum_{a\,|\,a\lessdot p,\ a\textup{ downward orphan}}w(a)+\sum_{b\,|\,b\gtrdot p}w(b)=2w(p)$ for all $p\in P$,
  3. (lem. 5.27(1)) the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$ for all $p\in P$, and
[…]
Proof.
This is proved by chaining together th. 5.2, lem. 5.7, lem. 5.17, lem. 5.18, lem. 5.23, and lem. 5.27. ∎
No upward or downward triple covers
Lemma 5.30.
No element of $P$ is covered by three or more distinct elements.
Proof.
The construction for this poof is shown in fig. 8.
Figure 8. Construction for the proof of lem. 5.30
Assume that we have $a\in P$ covered by distinct $p_{1}$, $q_{1}$ and $r_{1}$. By unique-cover-modularity, there is a $pq_{2}$ covering $p_{1}$ and $q_{1}$, a $qr_{2}$ covering $q_{1}$ and $r_{1}$, and an $rp_{2}$ covering $r_{1}$ and $p_{1}$. By lem. 5.25 (applied to $p_{1}$, $q_{1}$, and $r_{1}$), $pq_{2}$, $qr_{2}$, and $rp_{2}$ are distinct. By (26) (applied to $p_{1}$, $q_{1}$, and $r_{1}$), $w(a)=w(pq_{2})=w(qr_{2})=w(rp_{2})$.
By unique-cover-modularity, there is a $p_{3}$ covering $rp_{2}$ and $pq_{2}$, a $q_{3}$ covering $pq_{2}$ and $qr_{2}$, and an $r_{3}$ covering $qr_{3}$ and $rp_{3}$. Suppose two of $p_{3}$, $q_{3}$, and $r_{3}$ are the same, say $p_{3}$ is the same as $q_{3}$. Define $b=p_{3}=q_{3}$, so $b$ covers $pq_{2}$, $qr_{2}$, and $rp_{2}$. Then by (26) (applied to $pq_{2}$, $qr_{2}$, and $rp_{2}$), $w(b)=w(p_{1})=w(q_{1})=w(r_{1})$. Since $w$ is positive, (15) implies $2w(pq_{2})\geq w(p_{1})+w(q_{1})=2w(b)$. So $w(pq_{2})\geq w(b)$ and similarly $w(qr_{2})\geq w(b)$ and $w(rp_{2})\geq w(b)$. Applying (15) to $b$ and the above gives $2w(b)\geq w(pq_{2})+w(qr_{2})+w(rp_{2})\geq 3w(b)$. Since $w(b)>0$, this is a contradiction. Thus $p_{3}$, $q_{3}$, and $r_{3}$ are distinct.
By (26) (applied to $pq_{2}$, $qr_{2}$, and $rp_{2}$), $w(p_{3})=w(p_{1})$, $w(q_{3})=w(q_{1})$, and $w(r_{3})=w(r_{1})$. Applying (15) and positivity,
$2w(p_{1})+2w(q_{1})+2w(r_{1})$ $=2w(p_{3})+2w(q_{3})+2w(r_{3})$
$2w(p_{3})+2w(q_{3})+2w(r_{3})$ $\geq 2w(pq_{2})+2w(qr_{2})+2w(rp_{2})$ (27)
$2w(pq_{2})+2w(qr_{2})+2w(rp_{2})$ $\geq 2w(p_{1})+2w(q_{1})+2w(r_{1})$ (28)
Thus
$w(p_{1})+w(q_{1})+w(r_{1})$ $=w(pq_{2})+w(qr_{2})+w(rp_{2})=w(p_{3})+w(q_{3})+w(r_{3})$
and the inequalities (27) and (28) are equalities. Equation (15) and the assumption that $w$ is positive requires that $pq_{2}$, $qr_{2}$, $rp_{2}$, $p_{3}$, $q_{3}$, and $r_{3}$ cover no elements other than the ones mentioned above (and shown in fig. 8) and have no upward orphans. Applying (15) to $pq_{2}$, $qr_{2}$, and $rp_{2}$ thus gives
$w(a)=w(pq_{2})$ $=w(p_{1})/2+w(q_{1})/2$ (29)
$w(a)=w(qr_{2})$ $=w(q_{1})/2+w(r_{1})/2$ (30)
$w(a)=w(rp_{2})$ $=w(r_{1})/2+w(p_{1})/2.$ (31)
Together these prove $w(p_{1})=w(q_{1})=w(r_{1})=w(a)$ and $w(p_{3})=w(q_{3})=w(r_{3})=w(a)$.
By (16) and that $w$ is positive, $2w(a)\geq w(p_{1})+w(q_{1})+w(r_{1})=3w(a)$. Since $w(a)>0$, this is a contradiction.
Thus, there is no $a\in P$ that is covered by three distinct elements. ∎
Lemma 5.31.
No element of $P$ covers three or more distinct elements.
Proof.
This is proved dually to lem. 5.30, by interchanging the use of (15) and (16) and replacing “upward orphan” with “downward orphan”. ∎
Remark 5.32.
Lem. 5.30 and 5.31 are the deeper explanation of the long-known fact[Fom1994a]*sec. 2.2 that ${\operatorname{Ideal}_{\textup{f}}}(\mathbb{N}^{k})$ for $k>2$ has no positive weighting as a differential lattice. Similarly, they show why $\mathbb{SS}_{t}={\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{2t})={\operatorname{Ideal}_{\textup{f}}}(t\textit{-SkewStrip})$ is a Fomin lattice (para. 2 item (10)) but ${\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{kt})$ with $k>2$ has no positive weightings. (Compare $\mathbb{CP}_{32}=\mathcal{L}_{3,2}$ in [El2025a]*Fig. 2 with $\mathbb{CP}_{33}=\mathcal{L}_{3,3}$ in [El2025a]*Fig. 3.)
The equations implied by lem. 5.28 for any particular $p\in P$ are characterized by the structure of its neighborhood $N_{p}$ as a poset.
Lemma 5.33.
$N_{p}$ is characterized up to poset isomorphism fixing $p$ by the numbers $|C_{p}^{-}|$, $|C_{p}^{+}|$ and $|S_{p}|$: All elements of $C_{p}^{-}$ are covered by $p$, all elements of $C_{p}^{+}$ cover $p$, and $S_{p}$ forms a matching between subsets of $C_{p}^{-}$ and $C_{p}^{+}$, with each element of $S_{p}$ covering a distinct element of $C_{p}^{-}$ and being covered by a distinct element of $C_{p}^{+}$. In addition, $|C_{p}^{-}|\leq 2$, $|C_{p}^{+}|\leq 2$, $|S_{p}|\leq|C_{p}^{+}|$, $|S_{p}|\leq|C_{p}^{-}|$, and $|N_{p}|=|C_{p}^{-}|+|C_{p}^{+}|+|S_{p}|+1$.
Proof.
This follows from lem. 5.26, 5.27, 5.30, and 5.31. ∎
For each possible set of values of $|C_{p}^{-}|$, $|C_{p}^{+}|$, $|S_{p}|$, and $|N_{p}|$, fig. 5 shows the resulting structure of $N_{p}$, and the consequent constraints per lem. 5.28.
$|C_{p}^{+}|$ $|C_{p}^{-}|$ $|S_{p}|$ $|N_{p}|$ Neighborhood Constraints
0 0 0 1 Refer to caption $2w(p)=0$
0 1 0 2 Refer to caption $2w(p)=w(x_{1})$
0 2 0 3 Refer to caption $2w(p)=w(x_{1})+w(x_{2})$
1 0 0 2 Refer to caption $2w(p)=w(y_{1})$
1 1 0 3 Refer to caption $2w(p)=w(x_{1})+w(y_{1})$
1 1 1 4 Refer to caption $2w(p)=w(x_{1})$ $w(x_{1})=w(y_{1})$
1 2 0 4 Refer to caption $2w(p)=w(x_{1})+w(x_{2})+w(y_{1})$
1 2 1 5 Refer to caption $2w(p)=w(x_{1})+w(x_{2})$ $w(x_{1})=w(y_{1})$
Figure 9.1. The possible neighborhoods of $p$ and their constraints, part 1
$|C_{p}^{+}|$ $|C_{p}^{-}|$ $|S_{p}|$ $|N_{p}|$ Neighborhood Constraints
2 0 0 3 Refer to caption $2w(p)=w(y_{1})+w(y_{2})$
2 1 0 4 Refer to caption $2w(p)=w(x_{1})+w(y_{1})+w(y_{2})$
2 1 1 5 Refer to caption $2w(p)=w(x_{1})+w(y_{2})$ $w(x_{1})=w(y_{1})$
2 2 0 5 Refer to caption $2w(p)=w(x_{1})+w(x_{2})+w(y_{1})+w(y_{2})$
2 2 1 6 Refer to caption $2w(p)=w(x_{1})+w(x_{2})+w(y_{2})$ $w(x_{1})=w(y_{1})$
2 2 2 7 Refer to caption $2w(p)=w(x_{1})+w(x_{2})$ $w(x_{1})=w(y_{1})$ $w(x_{2})=w(y_{2})$
Figure 9.2. The possible neighborhoods of $p$ and their constraints, part 2
6. Global characterization
In this section, we continue by using the local characterization from sec. 5 to prove “global” facts about the poset $P$ of points. These culminate in an enumeration of a small number of possible structures for $P$.
In this section, we continue the assumptions we have made in section 5. Thus, all points $p\in P$ have a neighborhood $N_{p}$ as described by lem. 5.33, and $w$ must satisfy the consequent constraints per lem. 5.28.
Cartesian products
If a lattice is the cartesian product product of Fomin lattices, para. 2 item (12) shows that it is is a Fomin lattice. It is straightforward that the product lattice preserves any of the properties that we have been discussing (5.1, 5.13, and 5.24) that are true of the factors. Conversely, if a Fomin lattice is the cartesian product of nontrivial lattices, th. 2.8, lem. 2.9, 2.11, and 2.12 show that its factors are Fomin lattices that preserve any of the properties that are true of the product lattice. Thus, without excluding any Fomin lattices from our classification, we can restrict our attention to lattices that cannot be cartesian-factored.
Assumption 6.1.
Henceforth, we assume that $L$ cannot be factored into two nontrivial lattices, which is equivalent to assuming that $P$ cannot be partitioned into two disjoint non-empty posets $P=P_{1}\sqcup P_{2}$.
$P$ has a minimum element
Lemma 6.2.
Given any two elements $a,b\in P$, we can connect them by a sequence of elements
$a=x_{0},x_{1},x_{2},\ldots,x_{i},\ldots,x_{n-1},x_{n}=b$
with
$a=x_{0}\gtrdot x_{1}\gtrdot x_{2}\gtrdot\cdots\gtrdot x_{i}\lessdot\cdots\lessdot x_{n-1}\lessdot x_{n}=b,$
for some $0\leq i\leq n$. (That is, the sequence first descends, then ascends.)
Proof.
Because $P$ is locally finite and cannot be separated into the disjoint union of two non-empty posets, $a$ and $b$ can be connected with a sequence of elements $a=x_{0},x_{1},x_{2},\ldots,x_{n}=b$ where each adjacent pair $x_{i},x_{i+1}$ has either $x_{i}\lessdot x_{i+1}$ or $x_{i}\gtrdot x_{i+1}$. Given such a sequence, if there is an $i$ such that $0<i<n$ and $x_{i-1}\lessdot x_{i}\gtrdot x_{i+1}$, that is, $x_{i}$ is a local maximum, by the assumption of unique-cover-modularity, there is a $y$ such that $x_{i-1}\gtrdot y\lessdot x_{i+1}$. Thus $a=x_{0},x_{1},x_{2},\ldots,x_{i-1},y,x_{i+1},\ldots x_{n}=b$ is another such sequence.
We can iterate this transformation on the sequence. Each transformation preserves the number of adjacent pairs $x_{i}\gtrdot x_{i+1}$ and the number of adjacent pairs $x_{i}\lessdot x_{i+1}$, while moving the location of the $x_{i}\gtrdot x_{i+1}$ pairs earlier in the sequence. Thus any series of such transformations must terminate in a sequence for which there is no $i<i<n$ for which $x_{i}$ which is a local maximum. Thus, any terminal sequence must have some $0\leq i\leq n$ for which for all $0\leq j<i$, $x_{j}\gtrdot x_{j+1}$ and for all $i\leq j<n$, $x_{j}\lessdot x_{j+1}$. ∎
Lemma 6.3.
$P$ has a minimum element, which we denote $\hat{0}_{P}$.
Proof.
Since $P$ is non-empty and finitary by lem. 5.4(2), $P$ must contain at least one minimal element $m$. Assume there is an element $a\not\geq m$ in $P$. Then by lem. 6.2, there is a sequence of elements $m=x_{0},x_{1},x_{2},\ldots,x_{n}=a$ where for some $0\leq i\leq n$, for any $0\leq j<i$, $x_{j}\gtrdot x_{j+1}$ and for any $i\leq j<n$, $x_{j}\lessdot x_{j+1}$. But since $m$ is minimal, there is no $y$ with $m\gtrdot y$, and so $i=0$. If $n>0$, then $m=x_{0}\lessdot x_{1}\leq x_{n}=a$. That implies $m<a$, which contradicts that $a\not\geq m$. So $n=0$ and $m=x_{0}=a$, which also contradicts $a\not\geq m$. Thus there is no $a\not\geq m$ in $P$ and $m$ is the minimum of $P$. ∎
Lemma 6.4.
$P$ is graded. We define $\rho$ as the rank function on $P$. $\rho(\hat{0}_{P})=0$.
Proof.
This proof is a straightforward generalization of th. 2.3[Birk1967a]*§II.8 Th. 14, given that $P$ is locally finite, is unique-cover-modular, and has a minimum element $\hat{0}_{P}$. ∎
Remark 6.5.
In lem. 6.4, the hypothesis that $\hat{0}_{P}$ exists is necessary. A counterexample is the poset “3-Plait” in [Fom1994a]*Fig. 7 and Exam. 2.2.10. Further counterexamples are the posets of points of the lattices of cylindrical partitions.[El2025a]
The bottom elements of $P$
First we define the sequence, finite or infinite, of elements at the bottom of $P$ that form a chain, with each element (above $\hat{0}_{P}$) being the unique cover of the one below.
Definition 6.6.
We define $B_{1}=\hat{0}_{P}$. Then iteratively for each $i\geq 0$, if $B_{i}$ has a unique upward cover, we define $B_{i+1}$ to be that cover. If there is a $B_{\bullet}$ that does not have a unique cover (and thus by lem. 5.30 has either zero or two covers), we define $N$ to be the index of that $B_{\bullet}$. Otherwise, we define $N$ to be $\infty$. ($B$ for “bottom”.)
In the case $N=\infty$, the $B_{\bullet}$ form an upward semi-infinite path, and they are the entirety of $P$.
Figure 10. The case $N=\infty$ where the $B_{\bullet}$ form an upward semi-infinite path.
The constraints on $w$ are $w(B_{1})=r$, $2w(B_{1})=w(B_{2})$, and, for all $i\geq 2$, $2w(B_{i})=w(B_{i-1})+w(B_{i+1})$, which have the unique solution $w(B_{i})=ir$. Thus we have:
Case 6.7.
(para. 2 item (5)) $P$ is the upward semi-infinite path $B_{i}$ for $i\geq 1$ with $r>0$ and $w(B_{i})=ir$. $L$ is isomorphic to $\mathbb{N}$.
If $N$ is not $\infty$, then there are two cases: $B_{N}$ has zero upward covers and $B_{N}$ has two upward covers. If $B_{N}$ has zero upward covers, then $P$ consists of the finite path $B_{0}\lessdot B_{1}\lessdot B_{2}\lessdot\cdots\lessdot B_{N}$.
Figure 11. The case $N$ is not $\infty$, $B_{N}$ has zero upward covers, and the $B_{\bullet}$ form a finite path.
By lem. 5.28, the constraints on $w$ are
  • $w(B_{1})=r$;
  • if $N=1$, $2w(B_{1})=0$;
  • if $N\geq 2$, $2w(B_{1})=w(B_{2})$;
  • for all $2\leq i<N$, $2w(B_{i})=w(B_{i-1})+w(B_{i+1})$; and
  • if $N\geq 2$, $2w(B_{N})=w(B_{N-1})$ .
For all finite $N$, these imply $r=0$ and all $w(B_{i})=0$, so $w$ is not positive.
$P$ is two-dimensional
The remaining case is that $B_{N}$ has exactly two upward covers.
Definition 6.8.
We define $P^{+}$ as the elements of $P$ that are not a $B_{i}$ for which $i<N$.
Lemma 6.9.
$P^{+}$ has at least three elements: one is $B_{N}$ and two are the covers of $B_{N}$. All $p\in P^{+}$ are $\geq B_{N}$, and each has a finite grade $\rho(p)\geq\rho(B_{N})$ by lem. 6.4.
We will “draw the Hasse diagram of $P$ in the plane” by defining a set of locations $T_{xy}$ for all integers $x\geq 0$ and $y\geq 0$. As a poset, $T\cong\mathbb{N}\times\mathbb{N}$. Then we assign every $p\in P^{+}$ to a distinct location $T_{xy}$ with $x+y=\rho(p)-\rho(B_{N})=\rho(p)-N$. ($T$ for “top elements”.) That is, $B_{N}$ is assigned to $T_{00}$ and elements of $P^{+}$ are assigned “upward” from $T_{00}$. If $p\lessdot q$ in $P^{+}$ and $p$ is assigned to $T_{xy}$ then $q$ is assigned to either $T_{x+1,y}$ or $T_{x,y+1}$. Conversely, if both $T_{xy}$ and either $T_{x+1,y}$ or $T_{x,y+1}$ are assigned, the latter covers the former. Intuitively, the $T_{\bullet\bullet}$ form a quarter-plane extending upward, with $T_{00}$ at the bottom center, the $x$ coordinate increasing to the upper-left and the $y$ coordinate increasing to the upper-right. Note that all elements of $P^{+}$ are assigned to distinct locations $T_{\bullet\bullet}$, but some locations may not have an assigned element. We abuse notation by using $T_{xy}$ to denote both the location and, if an element of $P^{+}$ is assigned to that location, that element of $P^{+}$.
We define the coordinate $v=x+y$; $v$ is how far $T_{xy}$ is vertically above $T_{00}$. We define the coordinate h=yxh=y-x; $h$ is how far $T_{xy}$ is horizontally left of $T_{00}$ (if negative) or right of $T_{00}$ (if positive). Note that while $v$ ranges over all nonnegative integers, $h$ ranges over integers $v\leq h\leq v$ which have the same parity as $v$. The assignment of elements of $P^{+}$ realizes its Hasse diagram, with each cover relation being either one step to the upper left (incrementing $x$) or one step to the upper right (incrementing $y$). An example assignment is shown in fig. 12.
Figure 12. An example of assigning elements of $P^{+}$ to locations $T_{\bullet\bullet}$. Solid circles are assigned locations, empty circles are unassigned locations. Solid lines are coverings in $P$. Further elements are assigned to the upper-right of $T_{03}$, $T_{13}$, and $T_{23}$. $B_{i}$ for $i<N$ are assigned downward from $T_{00}$.
Figure 13. An example of assigning elements of $P^{+}$ with rank $v_{\textup{new}}+R$ to locations with $v=v_{\textup{new}}$ in steps (3a–f). $T_{ab}$ has $v=v_{\textup{new}}-1$ and $h=h^{\textup{min}}_{v_{\textup{new}}-1}$ and $T_{cd}$ has $v=v_{\textup{new}}-1$ and $h=h^{\textup{max}}_{v_{\textup{new}}-1}$.
We construct the assignments by first assigning $T_{00}=B_{N}$ and successively assigning each grade of $P^{+}$ above $N$ to locations so as to satisfy these rules.
Lemma 6.10.
We can assign the elements of $P^{+}$ to locations in $\{T_{xy}\,|\,x\geq 0\textup{ and }y\geq 0\}$ by the following process, which has the following properties:
  1. (1)
    Assign $B_{N}$ to $T_{00}$.
  2. (2)
    Assign rank N1N-1, the two covers of $B_{N}$, arbitrarily to $T_{10}$ and $T_{01}$.
  3. (3)
    Iteratively consider the grades of $P^{+}$ from $N+2$ upward. Terminate if we reach a grade that is empty (all elements of $P^{+}$ have been assigned):
    1. (a)
      Within grade $R$, we will assign the elements of that grade to distinct locations $T_{xy}$ with $x+y=v_{\textup{new}}=R-N$. At this point, these iteration hypotheses are true of the $T_{xy}$ with $x+y<v_{\textup{new}}$:
      1. (i)
        The elements assigned to locations with a particular value of $v<v_{\textup{new}}$ have values $h$ that are a contiguous subset of the integers $v\leq h\leq v$ which have the same parity as $v$. Define $h^{\textup{min}}_{v}$ to be the minimum $h$ of an assigned location with a particular value of $v$ and define $h^{\textup{max}}_{v}$ to be the maximum $h$ of an assigned location with a particular value of $v$. Similarly, define $x^{\textup{min}}_{v}$, $x^{\textup{max}}_{v}$, $y^{\textup{min}}_{v}$, and $y^{\textup{max}}_{v}$ as the minimum and maximum $x$ and $y$ coordinates of assigned locations with this value of $v$.
      2. (ii)
        For any $0<v<v_{\textup{new}}$, $h^{\textup{min}}_{v}=h^{\textup{min}}_{v-1}\pm 1$ and $h^{\textup{max}}_{v}=h^{\textup{max}}_{v-1}\pm 1$. Thus $x^{\textup{min}}_{v}=x^{\textup{min}}_{v-1}$ or $x^{\textup{min}}_{v-1}+1$, $x^{\textup{max}}_{v}=x^{\textup{max}}_{v-1}$ or $x^{\textup{max}}_{v-1}+1$, $y^{\textup{min}}_{v}=y^{\textup{min}}_{v-1}$ or $y^{\textup{min}}_{v-1}+1$, and $y^{\textup{max}}_{v}=y^{\textup{max}}_{v-1}$ or $y^{\textup{max}}_{v-1}+1$,
      3. (iii)
        If $p\lessdot q$ in $P^{+}$ are assigned at this point, then $q$ is assigned to some $T_{xy}$ and $p$ is assigned to either $T_{x-1,y}$ or $T_{x,y-1}$.
      4. (iv)
        Conversely, if $T_{xy}$ is assigned and either of $T_{x-1,y}$ or $T_{x,y-1}$ is assigned, then the former covers the latter in $P^{+}$.
    2. (b)
      Consider any adjacent pair of assigned locations with $v=v_{\textup{new}}-1$, $T_{xy}$ and $T_{x-1,y+1}$, with $x+y=v_{\textup{new}}-1$, $h^{\textup{min}}_{v_{\textup{new}}-1}\leq y-x$, and $(y+1)-(x-1)\leq h^{\textup{max}}_{v_{\textup{new}}-1}$. $T_{xy}$ and $T_{x-1,y+1}$ cover $T_{xy}\wedge T_{x-1,y+1}=T_{x-1,y}$ and so they must be covered by $T_{xy}\vee T_{x-1,y+1}$, which has rank $R$ in $P^{+}$. Assign $T_{xy}\vee T_{x-1,y+1}$ to location $T_{x,y+1}$. By lem. 5.31, $T_{xy}\vee T_{x-1,y+1}$ covers no element other than $T_{xy}$ and $T_{x-1,y+1}$.
    3. (c)
      Every element of rank $R$ not yet assigned necessarily covers at least one element of rank R1R-1. But any element of rank R1R-1 with $h^{\textup{min}}_{v_{\textup{new}}-1}<h<h^{\textup{max}}_{v_{\textup{new}}-1}$ already has two elements that cover it that were assigned in step (3b) that cover it, so the only elements of rank R1R-1 that it might cover are
      1. (i)
        the assigned location with minimum $h$, $T_{ab}=T_{x^{\textup{max}}_{v_{\textup{new}-1}},y^{\textup{min}}_{v_{\textup{new}-1}}}$ and/or
      2. (ii)
        the assigned location with maximum $h$, $T_{cd}=T_{x^{\textup{min}}_{v_{\textup{new}-1}},y^{\textup{max}}_{v_{\textup{new}-1}}}$
      (which might be the same location). However, it cannot cover both of these elements if they are distinct. If it did, then both of these elements would cover an element of rank R2R-2, which would have caused this element (of rank $R$) to have been assigned a location in step (3b).
    4. (d)
      If an unassigned element covers $T_{ab}$, but $T_{ab}$ has another assigned cover (which must be $T_{a,b+1}$), assign this element to $T_{a+1,b}$. There cannot be two such elements in a single rank, as then $T_{ab}$ would be covered by three elements.
    5. (e)
      If an unassigned element covers $T_{pq}$, but $T_{pq}$ has another assigned cover (which must be $T_{p+1,q}$), assign this element to $T_{p,q+1}$. There cannot be two such elements in a single rank, as then $T_{pq}$ would be covered by three elements.
    6. (f)
      The remaining case is $T_{ab}=T_{pq}$ and unassigned elements remain. By lem. 5.30, there are at most two such elements, which can be assigned arbitrarily to $T_{a+1,b}$ and $T_{a,b+1}$. (Ultimately, we will prove that this only happens in rank $N+1$, which was handled by step (2).)
    7. (g)
      At this point, all elements in rank $R$ of $P^{+}$ are assigned locations, and the iteration hypotheses (3(a)i)–(3(a)iv) are true of the $T_{xy}$ with $x+y\leq v_{\textup{new}}$.
The algorithm is nondeterministic in cases (2) and (3f); either cover of the unique element in one rank may be assigned to either of the two locations in the next rank. In all other steps, it is deterministic. The typical operation of steps (3b–f) is shown in fig. 13.
Proof.
The induction properties stated in the lemma and the correctness of the algorithm are proved by mutual induction. The details are tedious but straightforward. ∎
Because the assigned locations of each grade are contiguous and very similar to the assigned locations of the grades above and below, the possible $P^{+}$ are determined up to isomorphism by the outline of the set of assigned locations, the values $h^{\textup{min}}_{\bullet}$ and $h^{\textup{max}}_{\bullet}$.
The bottom of $P^{+}$
We can now examine the constraints implied by the neighborhoods of $P$.
Definition 6.11.
We define $w_{xy}=w(T_{xy})$ when $T_{xy}$ is allocated, $m=w(B_{N})=w_{00}$, $\delta^{-}=w_{10}-m$, and $\delta^{+}=w_{01}-m$. If $N>1$, we define $\epsilon=w_{B_{N-1}}-m$ and otherwise $\epsilon=-m$.
($m$ for the “weight of the middle element of $P^{+}$”, namely $T_{00}=B_{N}$; $\delta^{-}$ for “the difference in the negative direction of hh”, that is, toward the left side of $P^{+}$; $\delta^{+}$ for “the difference in the positive direction of hh”, that is, toward the right side of $P^{+}$.) Both $\delta^{-}$ and $\delta^{+}$ are well-defined since $T_{00}=B_{N}$ is assumed to have two upward covers, which are $T_{10}$ and $T_{01}$.
Lemma 6.12.
The constraints on $w$ due to applying lem. 5.28 to $B_{1}$ and all the $B_{i}$ for $i<N$ are
$w(B_{i})=ir\textup{\qquad\qquad for all }1\leq i\leq N$ (32)
In particular, $m=Nr$, $\epsilon=-r<0$, and $N=-m/\epsilon$.
Lemma 6.13.
The constraints on $w$ due to applying lem. 5.28 to $T_{\bullet 0}$ imply either
  1. (1)
    $\delta^{-}\geq 0$:we define $x_{L}$$\infty$, all $T_{\bullet 0}$ are allocated, and $w_{i0}=m+i\delta^{-}$ for all $0\leq i$; or
  2. (2)
    $\delta^{-}<0$:we define ${x_{L}}=-m/\delta^{-}$, $x_{L}$} is a positive integer $\geq 2$, $T_{i0}$ is allocated iff $i<{x_{L}}$, and $w_{i0}=m+i\delta^{-}$ for all $0\leq i<{x_{L}}$.
Proof.
This is proved by finding the smallest $I$ for which $T_{i0}$ is not allocated, if one exists. Because $T_{01}$ is allocated, unique-modular-covering guarantees that for any $T_{i0}$ with $i\leq I$ (any $i$ if $I$ does not exist), $T_{i1}$ is allocated and is not an upward orphan. The constraints on $w_{i0}$ are that $2w_{i0}=w_{i-1,0}+w_{i+1,0}$ for any $0<i<I$ and $2w_{i0}=w_{i-1,0}$ for $i=I$ (if $I$ exists). Thus, $w_{i0}=m+i\delta^{-}$ for any $0\leq i<I$ (and for any $0\leq i$ if $I$ does not exist). If no such $I$ exists, conclusion case (1) follows, since positivity for all of the $w_{\bullet 0}$ requires $\delta^{-}\geq 0$. If $I$ exists, conclusion case (2) follows, since the constraint on $w_{I-1,0}$ requires it to equal $\delta^{-}$. ∎
Lemma 6.14.
Similarly, the constraints on $w$ due to applying lem. 5.28 to $T_{0\bullet}$ imply either
  1. (1)
    $\delta^{+}\geq 0$:we define $y_{L}$$\infty$, all $T_{0\bullet}$ are allocated, and $w_{0i}=m+i\delta^{+}$ for all $0\leq i$; or
  2. (2)
    $\delta^{+}<0$:we define ${y_{L}}=-m/\delta^{+}$, $y_{L}$} is a positive integer $\geq 2$, $T_{0i}$ is allocated iff $i<{y_{L}}$, and $w_{0i}=m+i\delta^{+}$ for all $0\leq i<{y_{L}}$.
Definition 6.15.
We define $\Delta^{-}=-m/\delta^{-}$ and $\Delta^{+}=-m/\delta^{+}$, defining $m/0=\infty$. This parallels the relationship between $N$ and $\epsilon$, $N=-m/\epsilon$.
Lemma 6.16.
$\Delta^{-}$ is either $\infty$ (if $\delta^{-}=0$), a negative real number (if $\delta^{-}>0$), or the reciprocal of a positive integer (which is $x_{L}$}, if $\delta^{-}<0$). Similarly for $\Delta^{+}$ w.r.t. $y_{L}$}.
Lemma 6.17.
The constraint on $T_{00}$ is equivalent to
$\frac{1}{\Delta^{-}}+\frac{1}{\Delta^{+}}+\frac{1}{N}=1,$ (33)
defining $1/\infty=0$.
Proof.
Applying lem. 5.28 to $T_{00}$ yields $2w_{00}=(w_{00}+\epsilon)+w_{10}+w_{01}$, since $T_{10}$ and $T_{01}$ both exist and are upward orphans of $T_{00}$. Dividing this equation by m-m gives the conclusion. ∎
Lemma 6.18.
Not considering the order of the terms, simple enumeration shows the decompositions of $1$ as a sum of reciprocals as in (33) are:
$\frac{1}{1}$ $+\frac{1}{n}$ $+(-\frac{1}{n})\textup{\quad for any integer $n\geq 1$}$
$\frac{1}{1}$ $+\frac{1}{\infty}$ $+\frac{1}{\infty}$
$\frac{1}{2}$ $+\frac{1}{2}$ $+\frac{1}{\infty}$
$\frac{1}{2}$ $+\frac{1}{3}$ $+\frac{1}{6}$
$\frac{1}{2}$ $+\frac{1}{4}$ $+\frac{1}{4}$
$\frac{1}{3}$ $+\frac{1}{3}$ $+\frac{1}{3}$
The possible cases for $P$
Lemma 6.19.
Since
  1. (1)
    the possible decompositions of 1 as a sum of reciprocals are listed in lem. 6.18,
  2. (2)
    NN must be positive,
  3. (3)
    given the left-right symmetry of $P$, we conventionally choose ${x_{L}}\geq{y_{L}}$, and
  4. (4)
    $T_{10}$ and $T_{01}$ are allocated, so $\Delta^{-}\neq 1$ and $\Delta^{+}\neq 1$,
the possible cases for $P$ are limited to:
$\Delta^{-}$ $x_{L}$} $\Delta^{+}$ $y_{L}$} NN Lattice
n-n for $\geq 2 \infty nn$ $n$ 1 $\mathbb{Y}_{n}$
$\infty$ $\infty$ $\infty$ $\infty$ 1 $\mathbb{Y}$
$\infty$ $\infty$ 2 2 2 $\mathbb{SY}$
3 3 2 2 6 none (lem. 6.22
6 6 2 2 3 none (lem. 6.24
6 6 3 3 2 none (lem. 6.21)
4 4 2 2 4 none (lem. 6.23)
4 4 4 4 2 none (lem. 6.21)
3 3 3 3 3 none (lem. 6.21)
Remark 6.20.
If we do not require $\hat{0}$, then we remove the restriction $N>0$ in lem. 6.19. That allows the additional case $\Delta^{-}=\Delta^{+}=2$, and $N=\infty$. Developing that case leads to the Fomin lattices para. 2(11).
Lemma 6.21.
It is not possible that $x_{L}$} and $y_{L}$} are not $\infty$ and are >2>2.
Proof.
If $x_{L}$} and $y_{L}$} are not $\infty$ and are >2>2, then $T_{20}$ and $T_{02}$ are allocated and neither $T_{21}$ nor $T_{12}$ are upward orphans of $T_{11}$. Thus the constraints of lem. 5.28 for $T_{11}$ simplifies to $0=\delta^{-}+\delta^{+}$. But since $x_{L}$} and $y_{L}$} are finite, $\delta^{-}<0$ and $\delta^{+}<0$, which is a contradiction. ∎
Lemma 6.22.
It is not possible that ${x_{L}}=3$ and ${y_{L}}=2$.
Proof.
Figure 14. The case ${x_{L}}=3$ and ${y_{L}}=2$: Elements are labeled with their weights in multiples of $m$. Coverings marked with a stroke are upward orphan coverings. Double vertical strokes connect weights that are equal by lem. 5.27(2). Empty circles are $T_{\bullet\bullet}$ that are not allocated. The constraint on $T_{44}$ must be violated.
Applying the constraints of lem. 5.28 to $T_{\bullet\bullet}$ in increasing order of $v$ straightforwardly determines the allocations, non-allocations, and weights shown in fig. 14. But of necessity, $w_{44}=w_{33}=m$, which violates the constraint on $T_{44}$. ∎
Lemma 6.23.
It is not possible that ${x_{L}}=4$ and ${y_{L}}=2$.
Proof.
Figure 15. The case ${x_{L}}=4$ and ${y_{L}}=2$: Elements are labeled with their weights in multiples of $m$. Coverings marked with a stroke are upward orphan coverings. Double vertical strokes connect weights that are equal by lem. 5.27(2). Empty circles are $T_{\bullet\bullet}$ that are not allocated. The constraint on $T_{22}$ must be violated.
Applying the constraints of lem. 5.28 to $T_{\bullet\bullet}$ in increasing order of $v$ straightforwardly determines the allocations, non-allocations, and weights shown in fig. 15. But of necessity, $w_{22}=w_{11}=m$, which violates the constraint on $T_{22}$. ∎
Lemma 6.24.
It is not possible that ${x_{L}}=6$ and ${y_{L}}=2$.
Proof.
Figure 16. The case ${x_{L}}=6$ and ${y_{L}}=2$: Elements are labeled with their weights in multiples of $m$. Coverings marked with a stroke are upward orphan coverings. Double vertical strokes connect weights that are equal by lem. 5.27(2). Empty circles are $T_{\bullet\bullet}$ that are not allocated. The constraint on $T_{22}$ must be violated.
Applying the constraints of lem. 5.28 to $T_{\bullet\bullet}$ in increasing order of $v$ straightforwardly determines the allocations, non-allocations, and weights shown in fig. 16. But of necessity, $w_{22}=w_{11}=m$, which violates the constraint on $T_{22}$. ∎
Lemma 6.25.
For the first three cases of lem. 6.19, $L$ must be:
  1. (1)
    for $\Delta^{-}=-n$ and $\Delta^{+}=n$, $L$ is isomorphic to $\mathbb{Y}_{n}$ with weights that are $m/n$ times the canonical weights (para. 2(4) and fig. 3),
  2. (2)
    for $\Delta^{-}=\Delta^{+}=\infty$, $L$ is isomorphic to $\mathbb{Y}$ with weights that are $m$ times the canonical weights (para. 2(2) and fig. 1),
  3. (3)
    for $\Delta^{-}=\infty$ and $\Delta^{+}=2$, $L$ is isomorphic to $\mathbb{SY}$ with weights that are $m/2$ times the canonical weights (para. 2(3) and fig. 2),
with $P$ being the poset of join-irreducibles of $L$ with the corresponding weights.
Proof.
In all of these cases, the values $w_{B_{i}}$, $w_{x0}$, and $w_{0y}$ are determined (relative to $m$) by $x_{L}$}, $y_{L}$}, and $N$. Which other elements of $T_{\bullet\bullet}$ are allocated and their weights are straightforwardly determined by the constraints of lem. 5.28 on them (working in order of increasing $v$), and necessarily result in the known lattices listed in the table. ∎
7. Classification theorem
Theorem 7.1.
The only Fomin lattices $L$ that satisfy the assumptions
  1. (assump. 5.1) The lattice $L$ is distributive. We define the poset of points (join-irreducibles) to be $P$, so $L={\operatorname{Ideal}_{\textup{f}}}(P)$.
  2. (assump. 5.13) $P$ is unique-cover-modular.
  3. (assump. 5.24) The weighting $w$ is positive, that is, the set of values $V$, the target of $w$, is $\mathbb{Z}$, and all of the values of $w$ are positive. The differential degree $r$ is positive.
  4. (assump. 6.1) The lattice cannot be factored into two nontrivial lattices, which is equivalent to assuming that $P$ cannot be partitioned into two disjoint non-empty posets $P=P_{1}\sqcup P_{2}$.
are the cases
  1. $\mathbb{Y}$:
    (para. 2(2)) $P$ is isomorphic to the quadrant of the plane, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers $\geq 0$}\}$, and $L$ is isomorphic to $\mathbb{Y}$, Young’s lattice, the lattice of partitions, with weighting $r$ times the canonical weighting of 1.
  2. $\mathbb{SY}$:
[…]
  3. $\mathbb{Y}_{\bullet}$:
    (para. 2(4)) $P$ is isomorphic to the $k$-row strip for some $k\geq 1$, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers, $0\leq x$ and $0\leq y<k$}\}$, and $L$ is isomorphic to $\mathbb{Y}_{k}$, the $k$-row Young’s lattice, the lattice of partitions with at most $k$ parts, with weighting $r/k$ times the canonical weighting $w(T_{xy})=k+x-y$.
Proof.
Note that conclusion case $\mathbb{Y}_{k}$ comprises case 6.7, the upward semi-infinite path, as $\mathbb{Y}_{1}$ and the case lem. 6.25(1) as $\mathbb{Y}_{k}$ with $k\geq 2$. The theorem is the summation of case 6.7, lem. 6.19, 6.21, 6.22, 6.23, 6.24, and 6.25. ∎
8. Future directions
The strength of the assumptions of our classification theorem (th. 7.1 is disappointing, but they are satisfied by all known Fomin lattices except the one family that is not even distributive, the Young–Fibonacci lattices and cartesian products that include them. In future work, we hope to expand the classification to larger classes.
There are a number of ways of relaxing the assumption of positivity.
  1. (1)
    Considering sets of values $V$ that are not the integers $\mathbb{Z}$ may be of abstract interest but doesn’t seem to be directly relevant to Robinson–Schensted algorithms. However, investigations of “generic” weightings of a lattice can exploit this generality. For instance, looking at weightings of the quadrant as $P$, we could start with the generic formula $w((i,j))=\alpha+\beta r+\gamma i+\delta j$ and see what values of $\alpha$, $\beta$, $\gamma$, and $\delta$ allow the differential criterion to be satisfied. This can be modeled by setting $V=\mathbb{R}^{4}$ and translating $\alpha+\beta r+\gamma i+\delta j$ into $(\alpha,\beta,\gamma,\delta)$ and $r$ into $(0,1,0,0)$. Once the possible weightings for this $V$ are determined, then homomorphisms from $V$ to $\mathbb{Z}$ can be considered and the resulting weightings that are positive can be identified. From these positive weightings Robinson–Schensted algorithms can be constructed.
    When $L$, $w$, and $V$ are jointly constructed from some complex algebraic structure (such as [Fom1994a]*sec. 2.3) this sort of flexibility may be particularly useful.
  2. (2)
    Allowing values of $w$ to be negative (when $V=\mathbb{Z}$) does not seem to be relevant to constructing Robinson–Schensted algorithms.
  3. (3)
    Allowing values of $w$ to be 0 may be useful as an intermediate step in constructing positive weightings, but it will reveal no new Fomin lattices since th. 4.16 et al. show that a Fomin lattice with some edges weighted 0 can be truncated into connected components which are each Fomin lattices with all edges having nonzero weights.
  4. (4)
    Allowing $r\leq 0$ (with $V=\mathbb{Z}$) seems to have limited potential as then either $L$ is trivial or does not have $\hat{0}$. However, see the discussion of Fomin lattices without $\hat{0}$ in (5).
There are also a number of ways of relaxing our structural assumptions about $L$.
  1. (5)
    Relaxing the assumption that $L$ has a $\hat{0}$ seems like it would be tractable. But lem. 6.4 and rem. 6.5 show that the existence of $\hat{0}$ has non-obvious consequences in that it causes unique-cover-modularity to imply that $L$ is graded. Thus removing the requirement of $\hat{0}$ may introduce unexpected complexity.
    Currently, there is no known way to use a Fomin lattice without $\hat{0}$ to build a Fomin lattice with $\hat{0}$, and thus a Robinson–Schensted algorithm. But the truncation theorems (th. 4.16 et al.) suggest it’s worth looking for processes by which they might be exploited.
  2. (6)
    There are a considerable number of Fomin lattices which have all values of $w$ positive but $r=0$ (and necessarily, no $\hat{0}$): See para. 2(6)(8)(10)(11). There is no obvious application of these to Robinson–Schensted algorithms, as any cartesian product involving any of them has no $\hat{0}$ and cannot be used in Robinson–Schensted algorithms. But this class of lattices may have a different combinatorial significance.
  3. (7)
    Relaxing the assumption that poset of points is unique-cover-modular to just cover-modular seems like it may be tractable. The structure of neighborhoods would become more complicated but perhaps a careful analysis of the graphs of possible coverings would show that a tractable set of constraints characterize the weights of the points.
  4. (8)
    Relaxing the assumption that the poset of points is cover-modular seems to be more difficult as it is used in the proof that no element of $L$ has more than two upward or downward covers (lem. 5.30) and a number of the lemmas for global characterization.
  5. (9)
    Classifying modular but non-distributive Fomin lattices seems to be the most difficult problem. There seems to be no tractable representation of general modular lattices that parallels Birkhoff’s Representation Theorem for distributive lattices, and our entire analysis is based on characterizing $L$ in the manner of Birkhoff’s representation.
    On the other hand, the technique used in [Stan2012a]*Exer. 3.51(a) soln. can be restated to provide a way to analyze an arbitrary differential distributive lattice, or conversely, to provide a nondeterministic algorithm for constructing all differential distributive lattices for a wide range of definitions of “differential”. This can be used to prove that no point is covered by three distinct points (lem. 5.30) and that there is a unique minimum point (lem. 6.3). It may be possible to use this technique to prove additional lemmas used in this analysis without assuming unique-cover-modularity.
On the largest scale,
  1. (10)
    All of the Robinson–Schensted algorithms that have attracted attention in the past has been derivable as growth diagrams based on what we call Fomin lattices, viz., modular lattices whose weightings are constant on projectivity classes. However the growth diagram construction can be based on any pair of dual graded graphs which is a much broader and more complicated class of structures. There needs to be an investigation why only the Fomin lattices have appeared in previous works.
  2. (11)
    Richard Stanley asks how much of this analysis can be applied to sequentially differential posets[Stan1990a]*sec. 2141414Sequentially differential posets have a different differential degree for each grade of the poset. that are distributive lattices.
  3. (12)
    Jonathan Farley[Far2003a]*sec. 11 classifies all finitary distributive lattices which have a cover function, where there is a function $f:\mathbb{N}\rightarrow\mathbb{N}$ and each element of the lattice that covers $n$ elements is covered by $f(n)$ elements. This suggests extending his analysis can be extended to the weighted case, where
[…]